### v1

# A geometric determinant method and geometric dimension growth

Abstract.
We study a geometric version of the dimension growth conjecture. Let $X$ be an irreducible projective variety defined over ${\mathbb{C}}(t)$ of dimension $m$ and degree $d\geq 2$. The set $X(b)$ of ${\mathbb{C}}(t)$-rational points of degree smaller than $b$ lying on $X$, has a natural structure of an algebraic variety over ${\mathbb{C}}$. We study the dimension and irreducibility of this variety for each $b$, and obtain a geometric analogue of the classical dimension growth conjecture in this case, namely that $\dim X(b)\leq mb$ for any $d\geq 2$. In particular, if $X$ is defined over ${\mathbb{C}}$, this provides uniform bounds on the dimension of the space of degree $b$ rational curves lying on $X$.
By developing a geometric version of Heath-Brown’s $p$-adic determinant method for varieties defined over ${\mathbb{C}}(t)$, we are furthermore able to study and bound the number of irreducible components of largest possible dimension of $X(b)$, providing an upper bound polynomial in $d$, and independent of $b$, for $d\geq 6$. This also allows us to study $C(b)$ when $C$ is a projective or affine curve, and obtain an analogue of the Bombieri–-Pila theorem for affine curves (and similarly, a result for projective curves).
1. Introduction
Since the work of Bombieri–Pila [BP89], the determinant method has been a crucial tool in obtaining uniform upper bounds on integral and rational points on algebraic varieties. Through a $p$-adic adaptation of this determinant method, Heath-Brown [HB02] was able to prove the dimension growth conjecture for surfaces. This conjecture asserts more generally that an irreducible projective variety $X\subset{\mathbb{P}}^{n}$ of degree $d\geq 2$ over ${\mathbb{Q}}$ satisfies
$N(X,B)\ll_{n,d,\varepsilon}B^{\dim X+\varepsilon},$
where $N(X,B)$ counts the number of rational points of $X$ of height at most $B$. Note that the implicit constant in the inequality above depends only on $n,d$ and $\varepsilon$, and not on the variety $X$ itself. Using an affine variant of dimension growth, Browning–Heath-Brown–Salberger, and later Salberger were able to prove the dimension growth conjecture whenever $d\geq 4$ [BHBS06, Sal23b], while the case $d=2$ is known via different methods [HB02]. The case $d=3$ remains open, although a non-uniform version has been obtained by Salberger [Sal23b].
In recent years, people have become interested in understanding the implicit constant in dimension growth type bounds, and in particular in showing that it is polynomial in $d$. In addition, it was shown that these bounds can be taken to be independent of $\varepsilon$. Building upon work of various authors [Wal15, CCDN20, CDH+23, BCK25, BCN24] it is now known by work of Binyamini–Cluckers–Kato that if $X\subset{\mathbb{P}}^{n}$ is irreducible of degree $d\geq 4$ and defined over ${\mathbb{Q}}$, then
$N(X,B)\ll_{n}d^{2}B^{\dim X}\log(B)^{O(1)}.$
This bound is essentially optimal, in view of the recent work [CG25]. Let us also mention that nowadays dimension growth type bounds exist for all global fields, see [Sed17, CFL20, Ver22, PS22, CDH+23, Ver24].
For many counting problems over global fields, it is possible to formulate geometric or motivic analogues. If $X$ is a projective variety over ${\mathbb{C}}$, then for a positive integer $b$ the set
$\operatorname{Hom}_{b}({\mathbb{P}}^{1},X)=\{f:{\mathbb{P}}^{1}\to X\mid\deg f=b\}$
can be considered a suitable analogue for the set of rational points of (logarithmic) height $b$. Note that this set is an algebraic variety over ${\mathbb{C}}$, and hence the geometry of $\operatorname{Hom}_{b}({\mathbb{P}}^{1},X)$ is naturally of interest. For example, one can try to understand the dimension, irreducibility or singularities of $\operatorname{Hom}_{b}({\mathbb{P}}^{1},X)$, and these type of problems received significant attention in recent years [BV17, BB23, Fai22, LVW24, GHL24].
(1.1). Main results
The aim of this article is to continue this line of research into understanding the geometry of $X(b)$ for various $X$, and more specifically to study analogues of the dimension growth conjecture in these settings. Instead of working with varieties over ${\mathbb{C}}$, we work more generally with varieties over ${\mathbb{C}}(t)$. If $X\subset{\mathbb{P}}^{n}$ is a quasi-projective variety over ${\mathbb{C}}(t)$, then the set
$X(b)=\{(a_{0}:\ldots:a_{n})\in X({\mathbb{C}}(t))\mid a_{i}\in{\mathbb{C}}[t]\text{ and }\deg a_{i}<b\text{ for each }0\leq i\leq n\}$
has a natural structure of a quasi-projective variety over ${\mathbb{C}}$, and is the analogue of the set of rational points of height smaller than $b$ on $X$. Note that if $X$ is defined over ${\mathbb{C}}$ then $X(b)$ is the union of $\operatorname{Hom}_{i}({\mathbb{P}}^{1},X)$ for $i<b$. Our first main result is an analogue of dimension growth for the set $X(b)$.
Theorem A.
Let $X\subset{\mathbb{P}}^{n}$ be an irreducible projective variety of degree $d\geq 2$ and dimension $m$ defined over ${\mathbb{C}}(t)$. Then
$\dim X(b)\leq mb.$
Note that in this theorem any degree $d\geq 2$ is allowed, and in particular the case $d=3$. In addition, the above bound is attained as soon as $X$ contains a linear space of dimension $m-1$ defined over ${\mathbb{C}}$. One can ask for more geometric information about $X(b)$, such as its number of irreducible components. In fact, we believe that the natural analogue of dimension growth over ${\mathbb{Q}}$ should include some control on the number of irreducible components of $X(b)$ of largest possible dimension. Indeed, if a bound of the form
$N(X,B)\leq cB^{m}$
holds in the global field setting for some $c,m>0$, then the analogue over ${\mathbb{C}}(t)$ should be that the dimension of $X(b)$ is bounded by $mb$, and the number of irreducible components of $X(b)$ of dimension $mb$ is bounded by $c$. For such a result, we obtain the following.
Theorem B.
Let $X\subset{\mathbb{P}}^{n}$ be an irreducible projective variety of degree $d\geq 6$ and dimension $m$ defined over ${\mathbb{C}}(t)$. Then for $b\geq 1$ we have
$\dim X(b)\leq mb,$
and the number of irreducible components of $X(b)$ of dimension $mb$ is bounded by $O(d^{7})$. If $b\geq 7$ then this number of irreducible components is bounded by $O(d^{4})$.
The proofs of Theorems A and B both proceed via showing the stronger affine analogues Theorems 5.1 and 5.13, which are proven by induction on $m=\dim X$. For $X\subset{\mathbb{A}}^{n}$ a quasi-affine variety over ${\mathbb{C}}(t)$, we denote by
$X(b)=\{(a_{1},\ldots,a_{n})\in X({\mathbb{C}}(t))\mid a_{i}\in{\mathbb{C}}[t]\text{ and }\deg a_{i}<b\text{ for each }0\leq i\leq n\}$
the set of integral points of height at most $b$ on $X$. Again, $X(b)$ is naturally a quasi-affine variety over ${\mathbb{C}}$. A suitable uniform bound on $\dim X(b)$, with $X\subset{\mathbb{A}}^{n}$, yields both of the above theorems for projective varieties by passing to the affine cone. The reason for moving to the affine situation is that induction on dimension works much better, as the integral points of height at most $b$ in ${\mathbb{A}}^{n}$ are contained in a $b$-dimensional family of hyperplanes over ${\mathbb{C}}$.
The key difference between the proofs of Theorem A and B is the base of the induction.
For Theorem A we follow the approach outlined in [CDH+23], where the induction begins with $m=1$, i.e. the curve case. For this we prove a strong bound on the number of integral points of bounded height on plane affine curves with at least two points at infinity.
Proposition 1.1.
Let $C\subset{\mathbb{A}}^{2}$ be an irreducible curve over ${\mathbb{C}}(t)$. If $C$ has at least two distinct points at infinity, then for every $b\geq 1$ we have that
$\dim C(b)\leq 1.$
If $C$ has genus $1$, or genus 00 and at least $3$ points at infinity, Proposition 1.1 is closely related to Siegel’s theorem on integral points over function fields (see Theorem 2.13). If $C$ has genus at least $2$, then this theorem follows from the Mordell conjecture over function fields (see Theorem 2.14). Therefore the only remaining case is when $C$ is of genus 00 with $2$ points at infinity, for which analyzing Pell equations over function fields plays a role. Since it seems difficult to control the number of irreducible components of $C(b)$, this result does not suffice to prove Theorem B.
The proof of Theorem B is more involved. We follow the approach of [BHB06], where the base of the induction is for surfaces. To this aim, we develop a geometric version of the determinant method in the style of [BHB06] but with improvements from [CCDN20, Wal15]. Over ${\mathbb{Q}}$, the global determinant method of Salberger [Sal23a] yields the strongest results around dimension growth, in particular leading to a proof of uniform dimension growth for projective varieties of degree $d\geq 4$. This global determinant method constructs a single auxiliary variety catching all rational or integral points on the variety of interest. Such a construction seems impossible in the geometric setting, since an affine curve defined over ${\mathbb{C}}(t)$ can contain infinitely many integral points of bounded height. As such, it is not clear to us whether a geometric version of Salberger’s global determinant method exists. This is the reason why we restrict to $d\geq 6$ in Theorem B, while dimension growth over global fields is known for $d\geq 4$.
Remark 1.2.
One can approach Theorem A and Theorem B using spreading out methods [^1] We thank Tim Browning for pointing this out to us. as in [BV17]. The key point is to have good control of the dependence on $q$ for dimension growth results over ${\mathbb{F}}_{q}(t)$. For example, using [Ver24, Thm. 1.3], this approach leads to a quick proof of Theorem A when $X$ has degree $d\geq 65$. However, the more general results from [PS22, CDH+23, BCK25] for global fields leave the dependence on the field implicit, and while this dependence can likely be made explicit, this would require significant work. Additionally, we note that spreading out cannot be used to prove Theorem A when $d=3$, simply because dimension growth over ${\mathbb{F}}_{q}(t)$ remains unproven for $d=3$.
In addition to proving Theorem B, we use the determinant method to deduce an analogue of the Bombieri–Pila theorem for affine curves [BP89] and Heath–Brown’s bounds for projective curves [HB02]. While a geometric version of the Bombieri–Pila bounds already appears in [CCL15], where a bound on the dimension of $C(b)$ for an affine curve $C$ is given, bounding the number of irreducible components is new. The geometric version of Heath–Brown’s bound for projective curves is new in terms of both dimension and the number of irreducible components.
Theorem C (Affine curves).
Let $C\subset{\mathbb{A}}^{n}$ be an irreducible affine curve of degree $d$ over ${\mathbb{C}}(t)$. Then
$\dim C(b)\leq\left\lceil\frac{b}{d}\right\rceil.$
Moreover, the number of irreducible components of $C(b)$ of dimension $\lceil b/d\rceil$ is bounded by $O(d^{3}b)$.
Theorem D (Projective curves).
Let $C\subset{\mathbb{P}}^{n}$ be an irreducible projective curve of degree $d$ over ${\mathbb{C}}(t)$. Then
$\dim C(b)\leq\left\lfloor\frac{2(b-1)}{d}\right\rfloor+1\leq\left\lceil\frac{2b}{d}\right\rceil.$
Moreover, the number of irreducible components of $C(b)$ of dimension $\left\lfloor\frac{2(b-1)}{d}\right\rfloor+1$ is bounded by $O(d^{3}b)$.
These results follow from the more precise Theorem 4.3 and Proposition 4.4, which take into account the height of the defining polynomial of $C$.
For a projective surface one typically expects most rational points to accumulate on lines contained in the surface. Using the determinant method, we show that when removing all lines on our surface, one obtains strictly better bounds than in Theorem B. This is similar to classical dimension growth results in the literature (see e.g. [HB02, Theorem 10] or [Sal23b, Theorem 0.5]).
Theorem E (Projective surfaces).
Let $X\subset{\mathbb{P}}^{3}$ be an irreducible projective surface of degree $d\geq 6$ over ${\mathbb{C}}(t)$. Let $L$ denote the union of all lines $\ell\subset X$, and let $U\subset X$ be the complement of $L$ in $X$, which is a quasi-projective variety. Then
$\dim U(b)\leq\left(1+\max\left\{\frac{1}{2},\frac{3}{2\sqrt{d}}+\frac{1}{3}\right\}\right)b+2.$
Note that in the above result the quantity $\max\left\{\frac{1}{2},\frac{3}{2\sqrt{d}}+\frac{1}{3}\right\}$ is strictly smaller than $1$, since $d\geq 6$. Hence this result improves upon the bound coming from Theorem B (for $b\geq 5$).
Structure of the paper
In Section 2 we introduce needed definitions, and give some basic preliminary results such as a geometric version of the Schwartz–Zippel bounds and the Bombieri–Vaaler theorem. We also recall a version of Siegel’s theorem over function fields (due to Mason), and a version of the Mordell conjectures over function fields (due to Manin–Grauert). Finally, we provide some examples which illustrate the geometric behavior of $X(b)$ in several interesting situations. In Section 3, we prove Proposition 1.1, which forms the base of the induction for Theorem A. In Section 4 we develop a geometric version of the determinant method needed to control the number of irreducible components of largest possible dimension, and to prove Theorem B. We also prove Theorems C and D. In Section 5, we prove Theorems A, B and E. The projective case is deduced from the affine versions Theorems 5.13 and 5.1, which we prove by induction on dimension. The base case for this induction is counting on affine surfaces, which is given in Proposition 5.3.
Acknowledgements
The authors wish to thank Raf Cluckers for useful discussions, for suggesting the problem to them, and for his encouragement and support. The authors wish to thank Tim Browning for useful discussions. T. B. was supported by FWO Flanders (Belgium) with grant number 1131925N. Y. I. H. was partially supported by FWO Flanders (Belgium) with grant number 12B4X24N. F. V. was supported by FWO Flanders (Belgium) with grant number 11F1921N and by the Humboldt Foundation.
2. Preliminaries
(2.1). Definitions and notation
In general, throughout this paper we will denote $K={\mathbb{C}}(t)$ and ${\mathcal{O}}_{K}={\mathbb{C}}[t]$. Given $x=(x_{0}:\ldots:x_{n})$ in ${\mathbb{P}}^{n}(K)$, we may assume that the $x_{i}$ are in ${\mathcal{O}}_{K}$ and coprime. Then we define the height of $x$ to be
$h(x)=\max_{i}\deg x_{i}.$
Similarly, if $x=(x_{1},\ldots,x_{n})\in{\mathcal{O}}_{K}^{n}$ we define the height of $x$ to be
$h(x)=\max_{i}\deg x_{i}.$
For $X\subset{\mathbb{P}}^{n}_{K}$ a quasi-projective variety and $b$ a positive integer, we denote by
$X(b)=\{x\in X(K):h(x)<b\}\subset{\mathbb{P}}^{n}(K)$
the set of rational points of height at most $b$ on $X$. For a quasi-affine variety $X\subset{\mathbb{A}}^{n}_{K}$, we similarly define
$X(b)=\{x\in X\cap{\mathcal{O}}_{K}^{n}:h(x)<b\}\subset{\mathcal{O}}_{K}^{n}$
to be the set of integral points of height at most $b$ on $X$.
Lemma 2.1.
Let $b$ be a positive integer.
  1. (1)
    If $X\subset{\mathbb{A}}^{n}_{K}$ is a quasi-affine variety, then $X(b)$ is naturally a quasi-affine variety [^2] Note that $X(b)$ may not be reduced (see e.g. Example 2.18), in which case we can consider the associated reduced variety. For our purposes, this makes little difference. in ${\mathbb{A}}^{bn}_{{\mathbb{C}}}$.
  2. (2)
    If $X\subset{\mathbb{P}}^{n}_{K}$ is a quasi-projective variety, then $X(b)$ is naturally a quasi-projective variety in ${\mathbb{P}}^{b(n+1)-1}_{{\mathbb{C}}}$.
Proof.
For (1), suppose that $X$ is defined by the equations $f_{1}(x)=\ldots=f_{N}(x)=0,g(x)\neq 0$ with $f_{1},\ldots,f_{N},g\in{\mathbb{C}}[t][x_{1},\ldots,x_{n}]$. Taking $x_{i}=\sum_{j=0}^{b-1}a_{ij}t^{j}$ with $a_{ij}\in{\mathbb{C}}$, we obtain the equations for $X(b)$ by expanding every $f_{k}(x)$ into powers of $t$ and requiring that all coefficients coming from $f_{1},\ldots,f_{N}$ vanish as polynomials in $(a_{ij})_{ij}$, and at least one coefficient coming from $g$ does not vanish as a polynomial in $(a_{ij})_{ij}$.
For (2), we consider ${\mathbb{A}}^{b(n+1)}_{{\mathbb{C}}}$ as the space of $(n+1)$-tuples of polynomials of degree strictly smaller than $b$. Let $U\subset{\mathbb{A}}^{b(n+1)}_{{\mathbb{C}}}$ be the subset of such $(n+1)$-tuples which are coprime. Using non-vanishing of resultants, $U$ is an open and dense subset of ${\mathbb{A}}^{b(n+1)}$. It is also clear that $U$ is a cone with cone point 00. We can then define $X(b)$ as a subset of $U/{\mathbb{C}}^{\times}\subset{\mathbb{P}}^{b(n+1)-1}_{\mathbb{C}}$, and reason similarly as in the affine case. ∎
Given $X$ as above, we are interested in the geometric properties of $X(b)$. Evidently, if $X$ is defined over ${\mathbb{C}}$, then we see that $X(1)$ is isomorphic to $X({\mathbb{C}})$. Note that even if $X$ is smooth, $X(b)$ may be badly behaved (see Remark 2.20).
Definition 2.2.
Let $X\subset{\mathbb{P}}^{n}_{K}$ be a quasi-projective variety. We define
$N(X,b):=\dim X(b).$
If $X\subset{\mathbb{A}}^{n}_{K}$ is a quasi-affine variety, then we set
$N_{\mathrm{aff}}(X,b):=\dim X(b).$
The functions $N(X,b)$ and $N_{\mathrm{aff}}(X,b)$ are the analogues of the usual counting functions over global fields. Namely, they measure the density of ${\mathbb{C}}(t)$-points or ${\mathbb{C}}[t]$-points lying on $X$.
Remark 2.3.
Note that given a variety $X\subseteq{\mathbb{P}}^{n}_{{\mathbb{C}}(t)}$ and $b\geq 1$, the variety $X(b)$ depends on the embedding of $X$ into (${\mathbb{P}}^{n}_{{\mathbb{C}}(t)}$. For example, $C=\{yz^{d-1}=x^{d}\}\simeq{\mathbb{P}}^{1}_{{\mathbb{C}}(t)}$ but $C(b)$ is very different from (${\mathbb{P}}^{1}_{{\mathbb{C}}(t)}(b)$ (see Example 2.16).
(2.2). A trivial bound
We give a geometric version of the well known Schwartz–Zippel bounds, which include a bound on the number of irreducible components of largest possible dimension. The dimension bound is also proven in [CCL15, Lem. 5.1.1].
Proposition 2.4 (Geometric Schwartz–Zippel).
Let $X\subseteq{\mathbb{A}}^{n}_{K}$ be an affine variety of pure dimension $m$ and degree $d$. Then for every $b\geq 1$ we have
$N_{\mathrm{aff}}(X,b)\leq mb.$
Moreover, the number of irreducible components of $X(b)$ of dimension $mb$ is at most $d$.
Proof.
We may assume $X$ has strictly positive dimension, as the result is trivial otherwise. Let $Z$ be an irreducible component of $X$. Choose a coordinate $x_{j}$ such that $Z$ is not contained in $\{x_{j}=\lambda\}$ for any $\lambda\in K$, which is possible as $Z$ is irreducible of positive dimension. Consider the map $\pi_{j}:Z\to{\mathbb{A}}^{1}$ which projects onto the $j$-th coordinate. For every $y\in K$, the fibre $\pi_{j}^{-1}(y)$ is a variety of dimension $m-1$ and of degree at most $\deg Z$, if it is not empty. For each $b$, we get map $\pi_{j,b}:Z(b)\to{\mathbb{A}}^{1}_{K}(b)\simeq{\mathbb{A}}^{b}_{\mathbb{C}}$, where $\pi_{j,b}^{-1}(y)\simeq Z_{\tilde{y}}(b)$ for a suitable $\tilde{y}\in K$ of height smaller than $b$. By induction, for each $y\in{\mathbb{C}}^{b}$ we get $\dim\pi_{j,b}^{-1}(y)\leq(m-1)b$ with at most $\deg(Z)$ irreducible components of dimension $(m-1)b$. The claim now follows as
$N_{\mathrm{aff}}(Z,b)\leq b+(m-1)b=mb,$
and as the number of irreducible components of $Z(b)$ of dimension $mb$ is bounded by $\deg Z$ (note that $\pi_{j,b}$ is dominant when restricted to each irreducible component of dimension $mb$ of $Z(b)$ by dimension considerations). ∎
Note that Proposition 2.4 assumes no irreducibility of $X$, and hence the estimate is tight when $X$ is a union of linear spaces defined over ${\mathbb{C}}$. Moreover, this proposition also holds when $X$ is not defined over $K$.
Given a projective variety $X\subset{\mathbb{P}}_{K}^{n}$, we get the following projective version by applying Proposition 2.4 to the affine cone of $X$.
Corollary 2.5.
Let $X\subset{\mathbb{P}}^{n}_{K}$ be a projective variety of pure dimension $m$ and degree $d$. Then for every $b\geq 1$ we have
$N(X,b)\leq(m+1)b-1,$
and the number of irreducible components of $X(b)$ of dimension $(m+1)b-1$ is at most $d$.
Proof.
Let $Y\subset{\mathbb{A}}^{n+1}_{K}$ denote the affine cone of $X$. This is a variety of dimension $m+1$ and degree $d$, and hence we conclude by noting that
$N(X,b)=N_{\mathrm{aff}}(Y,b)-1$
and using Proposition 2.4. ∎
We deduce the following about existence of ${\mathbb{C}}$-points not lying on a given variety defined over $K$. This result is frequently used to show that given a generic condition over $K$, there already exists a ${\mathbb{C}}$-point satisfying the condition. See for example Proposition 5.10 or the proof of Lemma 5.6. This claim is also given in [CCL15, Corollary 5.1.2].
Corollary 2.6.
Let $X\subset{\mathbb{A}}^{n}_{K}$ be a quasi-affine variety with $\dim X<n$. Then there exists a point $y\in{\mathbb{C}}^{n}$ not belonging to $X(1)$. An analogous statement holds in the projective case.
Proof.
By Proposition 2.4, $\dim X(1)\leq\dim X$, whereas $\dim({\mathbb{A}}^{n}_{K}(1))=n>\dim X$, and so there exists a point $y\in{\mathbb{A}}^{n}_{K}(1)\setminus X(1)$. ∎
It is typically enough to deal with geometrically irreducible varieties.
Lemma 2.7.
Let $X$ be an affine or projective variety of dimension $m$ over $K$ which is integral over $K$, but not geometrically integral. Then there exists a subvariety $Y\subset X$ of dimension $m-1$, and of degree at most $d^{2}$, such that for all integers $b\geq 1$ we have
$X(b)\subset Y(b).$
In particular, if $X$ is projective then
$N(X,b)\leq mb-1,$
and the number of irreducible components of $X(b)$ of dimension $mb-1$ is at most $d^{2}$. If $X$ is affine then
$N_{\mathrm{aff}}(X,b)\leq(m-1)b,$
and the number of irreducible components of $X(b)$ of dimension $(m-1)b$ is at most $d^{2}$.
Proof.
This follows similarly to the argument below [HB02, Cor. 1]. The final part follows by applying Corollary 2.5 or Proposition 2.4 to $Y$. ∎
(2.3). Projections
To reduce to the case of hypersurfaces in our main results we use projections.
Proposition 2.8.
Let $X\subset{\mathbb{A}}^{n}_{K}$ be a geometrically integral affine variety of degree $d$ and dimension $m$. Then there exists a degree $d$ hypersurface $Z\subset{\mathbb{A}}_{K}^{m+1}$ birational to $X$, and such that for every $b\geq 1$ the varieties $X(b)$ and $Z(b)$ are birational. In particular
$N_{\operatorname{aff}}(X,b)=N_{\operatorname{aff}}(Z,b).$
Lemma 2.9.
Let $X\subset{\mathbb{P}}^{n}_{K}$ be a geometrically integral variety of dimension $m$ and degree $d\geq 2$ and let $L\subset{\mathbb{P}}^{n}_{K}$ be a hyperplane defined over ${\mathbb{C}}$ not containing $X$. Then there exists an $(m+1)$-dimensional coordinate space $\Gamma=\{x_{i_{1}}=\ldots=x_{i_{n-m-1}}=0\}\subset{\mathbb{P}}^{n}_{K}$ and a birational projection map $\pi:X\to Y$ to a degree $d$ geometrically integral hypersurface $Y\subset\Gamma$, such that for every $x\in X(K)$ we have,
$h(\pi(x))\leq h(x).$
In addition, the map $\pi$ can be chosen such that for every $x\in X$,
$x\in L\iff\pi(x)\in L.$
Proof.
We prove the claim by induction on $n$. If $X$ is a hypersurface there is nothing to prove, so assume $m\leq n-2$. For a choice of a generic point $p\notin X$, the projection of $X$ from $p$ to any hyperplane $\Gamma\subset{\mathbb{P}}^{n}_{K}$ not containing $p$ is birational to its image, which is a degree $d$ projective variety (see e.g. [BHB06, Page 8] or [Har95, Example 18.16]).
More precisely, by [Har95, Page 224], if we fix any $q$ on $X\backslash L$ then it is enough to choose $p$ not lying on the cone over $X$ with vertex $q$, which we denote by $C_{X,q}$. Note that $C_{X,q}$ does not contain $L$, as it is irreducible of dimension at most $n-1$, and contains $q\notin L$. Choose any $q\in X$, set $L_{i}=\{x_{i}=0\}$ and let $U$ be the complement of $C_{X,q}\cup L_{0}\cup\ldots\cup L_{n}$ in $L$ (if $L$ is a coordinate hyperplane, exclude $L$ from the union $L_{0}\cup\ldots\cup L_{n}$ above). Then $U$ is an open set in $L$, and by Corollary 2.6 and the above, we may find a point $p\in U$ defined over ${\mathbb{C}}$ and a coordinate hyperplane $L_{i}$ for some $1\leq i\leq n$, to obtain the desired projection $\pi:X\to Y\subset L_{i}$. Without loss of generality, assume $i=n$ and set $\Gamma=L_{n}$. Given $x\in X$, the image $\pi(x)$ is the intersection of $\Gamma$ with the line connecting $p$ and $x$ and is given by the formula
(1) $\pi(x)=\left(x_{0}-\frac{p_{0}}{p_{n}}x_{n},\ldots,x_{n-1}-\frac{p_{n-1}}{p_{n}}x_{n},0\right).$
Since $p$ is defined over ${\mathbb{C}}$, the inequality $h(\pi(x))\leq h(x)$ therefore holds. Finally, note that since $p\in L$, and $L$ contains any line between two points in $L$, we get
$\pi(x)\in L\iff x\in L.$
We conclude by induction, as $Y\subset\Gamma$ and $\dim\Gamma=n-1$. ∎
Proof of Proposition 2.8.
Consider the projective closure $\overline{X}$ of $X$ in ${\mathbb{P}}^{n}_{K}={\mathbb{A}}^{n}_{K}\cup L_{\infty}$, where $L_{\infty}=\{x_{0}=0\}$ is the hyperplane at infinity. By Lemma 2.9, there exists a coordinate linear space $\Gamma\subset{\mathbb{P}}^{n}_{K}$ of dimension $m+1$ and a regular birational map $\pi:\overline{X}\to Y$ to a degree $d$ hypersurface $Y\subset\Gamma$ which does not increase the height of each point $x\in\overline{X}(K)$. In addition, by the lemma we may choose $\pi$ such that $\pi(X)$ does not intersect the plane at infinity $L_{\infty}=\{x_{0}=0\}$ in ${\mathbb{P}}^{n}_{K}$. Without loss of generality, assume $\Gamma=\{x_{n}=0\}$. Set $Z:=\overline{\pi(X)}$. Then $Z$ is the required affine hypersurface. By Equation (1), since we project from points defined over ${\mathbb{C}}$, for each $b$ we get a birational projection map $\pi_{b}:X(b)\to Z(b)$. In particular, $N_{\operatorname{aff}}(X,b)=N_{\operatorname{aff}}(Z,b)$. ∎
(2.4). Linear spaces
In this section we give a precise count on points of bounded height on linear spaces, which will be used when counting on lines on a surface. We give a proof using geometry of numbers in function fields, although one can also prove this result via elementary means.
We denote by $\operatorname{Gr}(m,n)$ the Grassmanian of $m$-dimensional vector subspaces of $K^{n}$. For $L\in\operatorname{Gr}(m,n)$, we denote by $h(L)$ the height of the point $L$ under the Plücker embedding $\operatorname{Gr}(m,n)\to{\mathbb{P}}^{N}$. This quantity is also often called the determinant of $L$.
Lemma 2.10.
Let $L\in\operatorname{Gr}(m,n)$. Then for $b\geq h(L)$ we have that
$N_{\mathrm{aff}}(L,b)=bm-h(L).$
Moreover, for $b>h(L)/m$ we have that $L(b)\neq\varnothing$.
Proof.
Let $v_{1},\ldots,v_{m}\in{\mathcal{O}}_{K}^{n}$ be a reduced basis of $L$, i.e. ordered with increasing height such that each $v_{i}$ realizes the successive minimum $s_{i}=h(v_{i})$ of $L$ (see [Hes02, Lemma 4.1] or [Pau98]). Then it follows from [Hes02, Lem. 4.1] that
  1. (1)
    for $\lambda_{i}\in{\mathcal{O}}_{K}$ we have $h(\sum_{i=1}^{m}\lambda_{i}v_{i})=\max\limits_{i:\lambda_{i}\neq 0}\left(s_{i}+h(\lambda_{i})\right)$, and
  2. (2)
    $h(L)=s_{1}+\ldots+s_{m}$.
The first property shows that
$N_{\mathrm{aff}}(L,b)=\sum_{i=1}^{m}\max\{b-s_{i},0\}.$
Hence for $b\geq h(L)\geq s_{m}$, we obtain that $N_{\mathrm{aff}}(L,b)=bm-h(L)$. Moreover, for $b>h(L)/m\geq s_{1}$ we have that $N_{\mathrm{aff}}(L,b)>0$, as desired. ∎
Lemma 2.11.
Let $m<n$ and let $A\in K^{m\times n}$ be a matrix of full rank. Let $L\in\operatorname{Gr}(n-m,n)$ be the kernel of $A$. Consider $A$ as a point in $\operatorname{Gr}(m,n)$ by considering the rows of $A$ as vectors in $K^{n}$. We then have that
$h(L)=h(A).$
Proof.
After permuting the coordinates and Gaussian elimination, we may assume that $A$ is of the form
$A=\begin{pmatrix}I_{m}&\mid B\end{pmatrix}.$
The claim now follows as the columns of the following matrix form a basis of $L$
$\begin{pmatrix}-B^{T}\\ I_{m}\end{pmatrix}.\qed$
We will also need an analogue of the Bombieri–Vaaler theorem [BV83] for function fields, which follows from the above results.
Lemma 2.12 (Bombieri–Vaaler).
Let $A\in{\mathcal{O}}_{K}^{m\times n}$ be a matrix with $m<n$. Assume that $A$ has full rank. Then there exists a non-zero $x\in{\mathcal{O}}_{K}^{n}$ with $Ax=0$ and
$h(x)\leq\frac{h(\det(A\overline{A}^{T}))/2-h(D)}{n-m},$
where $D$ is the greatest common divisor of all $m\times m$ minors of $A$.
Proof.
Let $L\in\operatorname{Gr}(n-m,n)$ be the linear space corresponding to the kernel of $A$. Then Lemma 2.10 shows that there exists a nonzero point $x\in{\mathcal{O}}_{K}^{r}$ with $h(x)\leq h(L)/(n-m)$. The result then follows from Lemma 2.11 after noting that
$h(L)=h(A)=h(\det(A\overline{A}^{T}))/2-h(D).\qed$
(2.5). Points on curves over function fields
To prove Theorem A, we use induction on the dimension. The base for the induction is Proposition 1.1 about sparseness of ${\mathcal{O}}_{K}$-points lying on curves. For this aim, we need an analogue of Siegel’s theorem over function fields, which we recall from the literature below. The following is a result due to Mason.
Theorem 2.13 ([Mas84, Theorems 8, 11]).
Let $C\subset{\mathbb{A}}^{2}_{K}$ be a geometrically irreducible curve.
  1. (1)
    Assume that $C$ has genus 00 and at least three distinct points at infinity. Then $C$ has infinitely many ${\mathcal{O}}_{K}$-points if and only if there exists a rational parametrization $\psi:{\mathbb{A}}^{1}_{K}\to C$ of the form
    $\psi(x)=\left(\frac{a(x)}{c(x)},\frac{b(x)}{c(x)}\right)$
    where $a,b\in{\mathcal{O}}_{K}[x],c\in{\mathbb{C}}[x]$ have no common factors. Furthermore, if such a $\psi$ exists, then there are only finitely many ${\mathcal{O}}_{K}$-points on $C$ which are not the image under $\psi$ of a $K$-point on ${\mathbb{A}}^{1}_{K}$.
  2. (2)
    Assume that $C$ has genus $1$. Then $C$ has infinitely many ${\mathcal{O}}_{K}$-points if and only if there exists an elliptic curve $D\subset{\mathbb{A}}^{2}_{K}$ defined by $y^{2}=j(x)$ for a square-free monic polynomial $j\in{\mathbb{C}}[x]$ and a parametrization $\psi:D\to C$ of the form
    $\psi(x,y)=\left(\frac{a(x)+yb(x)}{c(x)},\frac{d(x)+ye(x)}{g(x)}\right),$
    where $a,b,d,e\in{\mathcal{O}}_{K}[x]$ and $c,g\in{\mathbb{C}}[x]$. Furthermore, in this case the ${\mathcal{O}}_{K}$-points of $C$ consist of the infinite family arising by taking $x\in{\mathbb{C}}$, and finitely many other ${\mathcal{O}}_{K}$-points.
To take care of the case of genus at least $2$, we recall the resolution of Mordell’s conjecture (Falting’s theorem) over function fields.
Theorem 2.14 (Manin–Grauert, see [Sam66]).
Let $C\subset{\mathbb{A}}^{2}_{K}$ be a geometrically irreducible curve of genus at least $2$ which is not defined over ${\mathbb{C}}$. Then $C({\mathcal{O}}_{K})$ is finite.
(2.6). Examples
In this section we provide examples which demonstrate some noteworthy phenomena arising in our geometric version of dimension growth.
Firstly, we show the upper bounds on dimension given in Theorems C and D are tight, where the curves attaining these bounds also attain analogous bounds in classical dimension growth.
Example 2.15 (See [CCL15, Remark 5.1.5]).
Let $d\geq 1$ and let $C\subset{\mathbb{A}}^{2}_{K}$ denote the affine curve defined by $y=x^{d}$. Then for every integer $b\geq 1$ we have
$N_{\mathrm{aff}}(C,b)=\left\lceil\frac{b}{d}\right\rceil,$
and the variety $C(b)$ is irreducible. Indeed, it is the image in ${\mathbb{A}}^{2}_{K}(b)$ of the irreducible variety ${\mathbb{A}}^{1}_{K}\left(\left\lceil\frac{b}{d}\right\rceil\right)$ under the map arising from $x\mapsto(x,x^{d})$. Note however that $C(b)$ is not reduced for every $b\geq 2$.
Example 2.16.
Let $d\geq 1$ and let $C\subset{\mathbb{P}}^{2}_{K}$ denote the projective curve defined by $yz^{d-1}=x^{d}$. Then for every integer $b\geq 1$ we have
$N(C,b)=2\left\lceil\frac{b}{d}\right\rceil-1,$
and the variety $C(b)$ is irreducible.
In Theorems C and D we only count the number of irreducible components of maximal possible dimension. Accessing lower-dimensional irreducible components seems quite difficult to us, and one should not expect any good bound on their number, as the following example illustrates.
Example 2.17.
For each $n\geq 1$, let $X_{n}$ be the vanishing locus of the generalized Pell equation $y^{2}-(t^{2}+t+1)x^{2}=(t-1)(t-2)...(t-n)$ in ${\mathbb{A}}^{2}_{K}$. Let $L$ be the function field $K(\sqrt{t^{2}+t+1})$ and let $\phi:L\to L$ be the non-trivial automorphism of $L$ over $K$.
By Lemma 3.2, $\dim X_{n}(b)=0$ for every $n,b$. Hence each $X_{n}(b)$ consists of a finite number of points. However, we claim that
#$\#X_{n}(b)\geq 2^{n},\quad\text{ for }b=n+1.$
Indeed, it is easy to see that for each $i\in{\mathbb{C}}$, the equation $y^{2}-(t^{2}+t+1)x^{2}=t-i$ has a solution $(x_{i},y_{i})\in{\mathcal{O}}_{K}^{2}$ with $\deg x_{i}=0$, $\deg y_{i}=1$. Now, for each function $\sigma:\{1,...,n\}\to\{1,-1\}$ we get an element
$y_{\sigma}+x_{\sigma}\sqrt{t^{2}+t+1}=\prod_{i=1}^{n}(y_{i}+\sigma(i)x_{i}\sqrt{t^{2}+t+1})\in L$
for which $(x_{\sigma},y_{\sigma})\in X_{n}(n+1)$.
We claim that these $(x_{\sigma},y_{\sigma})$ are all different. To see this, assume that $(x_{\sigma},y_{\sigma})=(x_{\tau},y_{\tau})$ for two functions $\sigma,\tau$. Let $S\subseteq\{1,\ldots,n\}$ be the set of $j$ for which $\sigma(j)\neq\tau(j)$ and define
$\alpha_{S}=\prod_{i\in S}(y_{i}+\sigma(i)x_{i}\sqrt{t^{2}+t+1})\in L.$
By assumption $\phi(\alpha_{S})=\alpha_{S}$. Therefore the norm of $\alpha_{S}$ is equal to $\alpha_{S}^{2}=\prod_{i\in S}(t-i)$. Each point $t-i$ for $i=1,\ldots,n$ is unramified in $L$, and so unique factorization into prime ideals implies that $S=\varnothing$. We conclude that #$\#X_{n}(n+1)\geq 2^{n}$, as desired.
Even if $X$ enjoys nice regularity properties, e.g. $X$ is a smooth variety, the scheme $X(b)$ may be very singular. In this respect, $X(b)$ is very different from the jet schemes of $X$ (see Remark 2.20).
Example 2.18.
Let $C$ denote the curve in ${\mathbb{P}}^{2}_{K}$ cut by $tx^{2}=yz$. Then $C(b)$ is non-reduced for every integer $b\geq 1$.
Proof.
Given $b\geq 1$, the scheme $C(b)$ is an open sub-scheme of the closed scheme $Y$ cut out in ${\mathbb{P}}^{3b-1}_{{\mathbb{C}}}$ by the equations obtained from the relation $tx^{2}=yz$. Setting $x=x_{0}+\ldots+x_{b-1}t^{b-1}$, we see that the homogeneous ideal ${\mathcal{I}}_{Y}$ defining $Y$ contains the element $x_{b-1}^{2}$, but since it is generated by homogeneous degree $2$ polynomials, $x_{b-1}\notin{\mathcal{I}}_{Y}$. It follows that $C(b)$ is non-reduced. ∎
Finally, given a variety $X$ and an integer $b\geq 1$, we are interested in the number of irreducible components of top dimension of $X(b)$. The next example shows that this number may strictly increase in $b$ even if $X$ is a smooth variety.
Example 2.19.
Let $X$ denote the affine surface in ${\mathbb{A}}^{3}_{K}$ cut out by the equation $xy=z$ and let $b\geq 1$ be an integer. Then $X(b)$ has dimension $b+1$ and $b$ irreducible components of top dimension.
Proof.
Consider the projection map $p_{b}:X(b)\to{\mathbb{A}}^{b}_{{\mathbb{C}}}$ defined by $p_{b}(x,y,z)=z$. If we fix $z\in{\mathcal{O}}_{K}$ non-zero, then up to multiplication by ${\mathbb{C}}^{\times}$ there are only finitely many $x,y\in{\mathcal{O}}_{K}$ such that $xy=z$. Hence every fibre of $p_{b}$ not over $z=0$ is one-dimensional (and the fibre over 00 is of dimension $b$). We therefore conclude that $\dim X(b)=b+1$. Now note that for each $10\leq i\leq b-$ 1, the subscheme
$X_{i}:=\{(x,y,z)\in X(b):x_{j}=y_{k}=0,\,j>i,k>b-1-i\}$
is irreducible of dimension $b+1$ since the projection $p_{i}:X_{i}\to{\mathbb{A}}^{i+1}\times{\mathbb{A}}^{b-i}$ by $p_{i}(x,y,z)=(x,y)$ is an isomorphism. ∎
Remark 2.20.
Let $X\subset{\mathbb{A}}^{n}_{K}$ be an affine variety defined over ${\mathbb{C}}$. Recall that the ${\mathbb{C}}$-points of the $k$-th jet scheme $J_{k}(X)$ of a scheme $X$ are given by $X({\mathbb{C}}[t]/(t^{k+1}))$, see e.g. [EM09] for more details. The space $X(b)$ is therefore naturally a subvariety of the $(b-1)$-st jet scheme $J_{b-1}(X)$ of $X$. While $J_{b-1}(X)$ reflects the singularities of the variety $X$, the variety $X(b)$ might be badly behaved and may have many irreducible components even if $X$ is smooth, as seen in Examples 2.18, 2.19. In contrast, if $X\subset{\mathbb{A}}^{n}_{K}$ is an irreducible local complete intersection with at most rational singularities, then $J_{b-1}(X)$ is irreducible of dimension $b\dim X$ by [Mus01].
3. Integral points on curves
In this section we prove Proposition 1.1, which will form the base of the induction for Theorem A.
We first need a result about Pell equations, for which we briefly deal with more general function fields. Let $L$ be a function field of a curve $C$ over ${\mathbb{C}}$, and let $S\subset C({\mathbb{C}})$ be a finite set of points. We will write the elements of $S$ as valuations on $L$. Denote by ${\mathcal{O}}_{L,S}$ the ring of $S$-integers of $L$, which consists of all elements $a\in L$ for which $v(a)\geq 0$ for every $v\not\in S$. For an element $a\in L$, recall that the height of $a$ is defined as
$h(a)=\deg(a),$
where $a$ is considered as a morphism $a:C\to{\mathbb{P}}^{1}$. When $L=K$ and $S=\{\infty\}$, this height agrees with the height as defined in the previous section.
Lemma 3.1.
Let $b$ be a positive integer. Then the set
${\mathcal{O}}_{L,S}^{\times}(b)=\{a\in{\mathcal{O}}_{L,S}^{\times}:h(a)<b\}$
is finite up to multiplication by ${\mathbb{C}}^{\times}$.
Proof.
Let $S=\{s_{1},\ldots,s_{n}\}$ and let $a$ be an element of ${\mathcal{O}}_{L,S}^{\times}(b)$. Then all zeroes and poles of $a$ are contained in $S$. Denoting by $e_{i}$ the order of $a$ at $s_{i}$, we obtain that
$\deg a=\sum_{e_{i}\geq 0}e_{i}=\sum_{e_{i}<0}-e_{i}<b.$
Hence there are only finitely many options for $e_{1},\ldots,e_{n}$. We conclude by noting that an element of $L$ is determined by its zeroes and poles with multiplicities, up to multiplication by ${\mathbb{C}}^{\times}$. ∎
We need some auxiliary results about equations of Pell type to deduce Proposition 1.1.
Lemma 3.2 (Pell equation).
Let $\beta,\gamma$ be in ${\mathcal{O}}_{K}$ and let $b$ be a positive integer. Assume that $\beta$ is not a square in $K$ and that $\gamma$ is non-zero. Then the equation
$x^{2}-\beta y^{2}=\gamma$
has only finitely many solutions $a\in{\mathcal{O}}_{K}^{2}$ with $h(a)\leq b$.
Proof.
Let $L=K(\sqrt{\beta})$, which is the function field of an irreducible curve $D$ since $\beta$ is not a square in $K$. The curve $D$ comes with a degree two cover $D\to C$. Let $S\subset D$ consist of all points above $\infty$ and all points above zeroes of $\gamma$. Let $(x_{1},y_{1})$ and $(x_{2},y_{2})$ be two solutions of $x^{2}-\beta y^{2}=\gamma$ in ${\mathcal{O}}_{K}^{2}$ of height at most $b$. Then the elements $x_{i}+y_{i}\sqrt{\beta}$ in ${\mathcal{O}}_{L,S}$ have height at most $2b+h(\sqrt{\beta})$, and so the element $z=(x_{1}+y_{1}\sqrt{\beta})/(x_{2}+y_{2}\sqrt{\beta})$ has height bounded by $b^{\prime}=4b+2h(\sqrt{\beta})$. But $z$ is also a unit in ${\mathcal{O}}_{L,S}$ and so $z\in{\mathcal{O}}_{L,S}^{\times}(b^{\prime})$. We conclude by the previous lemma, and the fact that the equation $x^{2}-\beta y^{2}=\gamma$ is non-homogeneous (otherwise the set of solutions is a cone). ∎
To prove Proposition 1.1, we will distinguish several cases. Assume first that $C$ has genus at least one. If $C$ is not defined over ${\mathbb{C}}$, then Siegel’s theorem over function fields gives finiteness of $C({\mathcal{O}}_{K})$ (Theorems 2.13(2) and 2.14). If $C$ is defined over ${\mathbb{C}}$, then the following geometric argument shows that $C(K)=C({\mathbb{C}})$.
Lemma 3.3.
Let $C\subseteq{\mathbb{A}}^{2}_{\mathbb{C}}$ be a geometrically integral curve defined over ${\mathbb{C}}$ of genus at least $1$. Then $C(K)=C({\mathbb{C}})$. In particular, $\dim C(b)\leq 1$ for every positive integer $b$.
Proof.
Let $C_{s}\to C$ be the normalization of $C$, which is a smooth projective curve birational to $C$. Any $K$-point of $C$ corresponds to a morphism ${\mathbb{P}}_{\mathbb{C}}^{1}\to C$, which yields a map ${\mathbb{P}}_{\mathbb{C}}^{1}\to C_{s}$ since ${\mathbb{P}}^{1}$ is normal. As $C_{s}$ has genus at least $1$, this morphism must be constant, i.e. it must correspond to a ${\mathbb{C}}$-point on $C$.
For the final statement simply note that every ${\mathcal{O}}_{K}$-point is also a $K$-point. ∎
For the genus 00 case, we will need the following classical lemma to parametrize $C$. We embed ${\mathbb{A}}^{2}_{K}$ in ${\mathbb{P}}^{2}_{K}$ so that the line at infinity is given by $x_{2}=0$.
Lemma 3.4.
Let $C\subset\mathbb{P}_{K}^{2}$ be a geometrically integral projective curve over $K$. Suppose that $C$ has genus 00 and that $C(K)$ has at least three smooth points. Then there exist homogeneous polynomials $P_{1},P_{2},P_{3}\in{\mathcal{O}}_{K}[s,t]$ all of the same degree such that
$\mathbb{P}^{1}_{K}\to C:[s:t]\mapsto[P_{1}(s,t):P_{2}(s,t):P_{3}(s,t)]$
is a birational map. Moreover, if $C$ has at least three points at infinity and infinitely many points of the form $[x:y:1]$, with $x,y\in{\mathcal{O}}_{K}$, then we can take $P_{3}$ to be a polynomial over ${\mathbb{C}}$.
Proof.
Let $\overline{K}$ be the algebraic closure of $K$ and let $C_{s}\to C$ be the normalization of $C$. Then $C_{s,\overline{K}}$ is an irreducible smooth, projective, genus 00 curve and hence isomorphic to $\mathbb{P}_{\overline{K}}^{1}$. Therefore we can find an isomorphism $\psi_{1}:\mathbb{P}_{\overline{K}}^{1}\to C_{s,\overline{K}}$, which leads to a birational map
$\psi:\mathbb{P}_{\overline{K}}^{1}\to C_{\overline{K}}:[s:t]\mapsto[P_{1}(s,t):P_{2}(s,t):P_{3}(s,t)],$
where the $P_{i}$ are homogeneous polynomials over $\overline{K}$. Moreover, by composing with an automorphism of $\mathbb{P}_{\overline{K}}^{1}$ we may assume that $\psi([0:1])$, $\psi([1:0])$ and $\psi([1:1])$ are smooth $K$-points on $C(K)$. We may also assume that $P_{3}$ has $1$ as one of its coefficients.
We now claim that $P_{i}$ have coefficients $K$. To see this, consider any $\sigma\in\text{Gal}(\overline{K}/K)$. Let $\psi^{\sigma}:\mathbb{P}_{\overline{K}}^{1}\to\mathbb{P}_{\overline{K}}^{2}$ be the map $\psi$ with all the coefficients of the $P_{i}$ replaced by their images under $\sigma$. Note that $\psi^{\sigma}\left(\mathbb{P}_{K}^{1}\right)=\sigma\left(\psi\left(\mathbb{P}_{K}^{1}\right)\right)$. As $C$ is defined over $K$, the latter is actually a subset of $C_{\overline{K}}$. As a consequence, the image of $\psi^{\sigma}$ is also $C_{\overline{K}}$.
Now note that $\psi^{\sigma}$ pulls back to a map $\psi^{\sigma}_{1}:\mathbb{P}_{\overline{K}}^{1}\to C_{s,\overline{K}}$. Moreover, because $\psi^{\sigma}$ is injective on an open subset of $\mathbb{P}_{K}^{1}$, $\psi^{\sigma}_{1}$ has degree $1$ and is therefore an isomorphism. Hence $\psi_{1}^{\sigma}=\psi_{1}\circ\alpha$ for some automorphism $\alpha$ of $\mathbb{P}_{K}^{1}$. Because both $\psi$ and $\psi^{\sigma}$ take the same values on $[0:1]$, $[1:0]$ and $[1:1]$, it follows that $\alpha$ is the identity, so $\psi=\psi^{\sigma}$. As $P_{3}$ had $1$ as one of its coefficients, it follows that $P_{i}^{\sigma}=P_{i}$. We conclude $P_{i}$ are defined over $K$. The final claim follows from Theorem 2.13. ∎
Proof of Proposition 1.1.
Assume first that $C$ has genus at least $1$. If $C$ is defined over ${\mathbb{C}}$, then the result follows from Lemma 3.3. So assume that $C$ is not defined over ${\mathbb{C}}$. If $C$ has genus at least $2$ then the result follows from Theorem 2.14. If $C$ has genus $1$, then by Mason’s theorem 2.13, the ${\mathcal{O}}_{K}$-points consist of an infinite family parameterized by ${\mathbb{C}}$, and finitely many other ones, so we get $\dim C(b)\leq 1$ as well.
So now assume that $C$ has genus 00. By Lemma 3.4, there exist homogeneous polynomials $P_{1},P_{2},P_{3}\in{\mathcal{O}}_{K}[s,t]$ such that all but finitely many integral points on $C$ are of the form
$\left(\frac{P_{1}(u,v)}{P_{3}(u,v)},\frac{P_{2}(u,v)}{P_{3}(u,v)}\right).$
Now we need to find $[u:v]\in{\mathbb{P}}_{K}^{2}$ for which $P_{1}(u,v)/P_{3}(u,v)$ and $P_{2}(u,v)/P_{3}(u,v)$ are both elements of ${\mathcal{O}}_{K}$. We may rewrite these rational functions as $Q_{1}/R_{1}$, $Q_{2}/R_{2}$, where $Q_{i}\in{\mathcal{O}}_{K}[u,v]$ and $R_{i}\in{\mathcal{O}}_{K}[u,v]$ have no common factors over $K[u,v]$. By taking the homogeneous resultant, we may find $A_{i},B_{i},C_{i},D_{i}\in{\mathcal{O}}_{K}[u,v]$ with
$A_{i}Q_{i}+B_{i}R_{i}=x_{i}u^{k_{i}}$
and
$C_{i}Q_{i}+D_{i}R_{i}=y_{i}v^{\ell_{i}},$
where $x_{i},y_{i}\in{\mathcal{O}}_{K}$ and $k_{i},\ell_{i}\in{\mathbb{N}}$. Therefore, we see that if $Q_{i}(u,v)/R_{i}(u,v)\in{\mathcal{O}}_{K}$ for a certain choice of $(u,v)\in{\mathcal{O}}_{K}^{2}$, then $R_{i}(u,v)\mid x_{i}u^{k_{i}}$ and $R_{i}(u,v)\mid y_{i}v^{\ell_{i}}$. Since we may assume that $u$ and $v$ have no common factors, we conclude that $R_{i}(u,v)\mid x_{i},y_{i}$, proving that $R_{i}$ can only take finitely many values, up to a unit (i.e. an element of ${\mathbb{C}}^{\times}$), as any element of ${\mathcal{O}}_{K}$ only has finitely many divisors.
Assume now that $C$ has at least three distinct points at infinity. Then we moreover have that $P_{3}$ and hence $R_{1},R_{2}$ are polynomials over ${\mathbb{C}}$. Therefore, both of $R_{1},R_{2}$ factor over ${\mathbb{C}}$ into linear factors. As there are at least two different points at infinity, at least two among these factors must be linearly independent, say $au+bv$ and $a^{\prime}u+b^{\prime}v$. Now, because $R_{1}$ and $R_{2}$ can take only finitely many values up to a unit, the same is true for $a^{\prime}u+b^{\prime}v$ and $au+bv$. Moreover, because the choice of $[u:v]$ is homogeneous in $u,v$, we may assume that $au+bv$ takes only finitely many values (by dividing out in $u,v$ by any potential unit). Then $a^{\prime}u+b^{\prime}v$ takes values in a one-dimensional space, and therefore so do $u$, $v$, showing that indeed $\dim C(b)\leq 1$.
Finally, if $C$ contains precisely two points at infinity, then the $R_{i}$ must be either of the form $L_{2}^{r_{i}}$ for $L_{2}\in{\mathcal{O}}_{K}[s,t]$ of degree $2$ over $K$, or of the form $L_{1}^{r_{1}}L_{2}^{r_{2}}$ for two linearly independent linear forms. In the former case, $L_{2}$ can take only finitely many values, and we conclude by Lemma 3.2 (by completing the square); in the latter case, we can use precisely the same strategy as before. ∎
4. The determinant method
Our main technique for proving Theorem B is a geometric analogue of the determinant method. This method allows one to construct auxiliary hypersurfaces capturing rational points of bounded height lying on $X$, and is the main topic of this section. Over global fields, this method has a long history with many improvements over the years. Let us mention the work of Bombieri–Pila [BP89] where the determinant method first appears, Heath–Brown’s $p$-adic adaptation [HB02], Salberger’s global improvement using many primes [Sal23b], Walsh’ work to keep track of the height of the polynomial [Wal15] and finally a version which is polynomial in $d$ due to Castryck–Cluckers–Dittmann–Nguyen [CCDN20].
Our geometric determinant method is a mixture of these works. It resembles Heath–Brown’s $p$-adic determinant method for multiple primes at the same time, but with the improvements by Walsh and Castryck–Cluckers–Dittmann–Nguyen.
We also prove Theorems C and D in this section.
(4.1). A determinant lemma
We need a lemma on divisibility of certain determinants.
For $X\subset{\mathbb{P}}^{n+1}_{{\mathcal{O}}_{K}}$ and $p\in{\mathcal{O}}_{K}$ a non-zero prime element, we denote by $X_{p}\subset{\mathbb{P}}^{n+1}_{\mathbb{C}}$ the reduction of $X$ modulo $p$. If $X$ is described by a set of polynomial equations over ${\mathcal{O}}_{K}$ and $p=t-\lambda\in{\mathcal{O}}_{K}$, then $X_{p}$ is obtained by simply substituting $t=\lambda$ in the equations for $X$.
Lemma 4.1 (Divisibility of determinants).
Let $X\subset{\mathbb{P}}^{n+1}_{{\mathcal{O}}_{K}}$ be a geometrically integral hypersurface, and for a non-zero prime element $p\in{\mathcal{O}}_{K}$ let $X_{p}$ be the reduction of $X$ modulo $p$. Let $P$ be a ${\mathbb{C}}$-point on $X_{p}$ of multiplicity $\mu$, and let $a_{1},\ldots,a_{s}$ be ${\mathcal{O}}_{K}$-points on $X$, given by primitive ${\mathcal{O}}_{K}$-coordinates, that all reduce to $P$ modulo $p$. Let $F_{1},\ldots,F_{s}$ be homogeneous polynomials in $x_{0},\ldots,x_{n+1}$ over ${\mathcal{O}}_{K}$. Then $\det(F_{i}(a_{j}))_{i,j}$ is divisible by $p^{e}$ with
$e\geq\left(\frac{n!}{\mu}\right)^{1/n}\frac{n}{n+1}s^{1+1/n}-O_{n}(s).$
Proof.
This follows exactly as in [Sal07, Lem. 2.5]. To make the upper bound independent of $d$, one follows [CCDN20, Cor. 2.5]. ∎
(4.2). Auxiliary polynomials
For the construction of our auxiliary polynomials, we need a coordinate transformation which keeps track of the height of our equations. For a homogeneous polynomial $f\in{\mathcal{O}}_{K}[x_{0},\ldots,x_{n+1}]$ denote by $c_{f}$ the coefficient of $x_{n+1}^{d}$ in $f$, and by $h(f)$ the height of the polynomial $f$ considered as a point in projective space ${\mathbb{P}}^{N}(K)$. The following is clear.
Lemma 4.2.
Let $f\in{\mathcal{O}}_{K}[x_{0},\ldots,x_{n+1}]$ be primitive and homogeneous of degree $d$ and set $X=V(f)$. Take $a_{0},\ldots,a_{n}\in{\mathbb{C}}$ such that
$h(f(a_{0},\ldots,a_{n},1))=h(f),$
and consider the invertible linear transformation
$A(x_{0},\ldots,x_{n+1})=(x_{0}+a_{0}x_{n+1},\ldots,x_{n}+a_{n}x_{n+1},x_{n+1}).$
Then $g(x):=f(A(x))$ is a primitive and homogeneous polynomial defined over ${\mathcal{O}}_{K}$ of degree $d$ such that $h(g)=h(c_{g})=h(f)$. Moreover, denoting $Y=V(g)$, the varieties $X(b)$ and $Y(b)$ are isomorphic for every integer $b\geq 1$.
Theorem 4.3 (Projective auxiliary polynomials).
Let $f\in{\mathcal{O}}_{K}[x_{0},\ldots,x_{n+1}]$ be absolutely irreducible, primitive and homogeneous of degree $d$, let $X$ be the variety cut out by $f$ in ${\mathbb{P}}^{n+1}_{K}$, and let $b$ be a positive integer. Take positive integers $\mu_{1},\ldots,\mu_{\ell}$ so that
$\sum_{i=1}^{\ell}\frac{1}{\mu_{i}^{1/n}}\geq\frac{(n+1)(b-1)}{nd^{1/n}}-\frac{h(f)-1}{nd^{1+1/n}},$
and let $p_{1},\ldots,p_{\ell}\in{\mathcal{O}}_{K}$ be non-zero primes. Then if $P=(P_{1},\ldots,P_{\ell})\in\prod X_{p_{i}}({\mathbb{C}})$ is a tuple where each $P_{i}$ is of multiplicity $\mu_{i}$, then there exists a polynomial
$g:=g_{P}\in{\mathcal{O}}_{K}[x_{0},\ldots,x_{n+1}]$
coprime to $f$, of degree $O_{n}(d^{1+1/n}b+d^{3}\ell)$, and vanishing on
$X(b;P_{1},\ldots,P_{\ell})=\{x\in X(b):x\equiv P_{i}\bmod p_{i}\}.$
The most important use of this theorem is when all $\mu_{i}$ are equal to $1$. In that case, one simply takes $\ell$ primes $p_{1},\ldots,p_{\ell}$ and smooth points $P_{i}\in X_{p_{i}}({\mathbb{C}})$. An application of this case is enough to deduce Theorem D. However, to obtain Theorem B for small values of $d$, namely $6\leq d\leq 9$, we will need to apply this result with some singular points on the reductions $X_{p_{i}}$.
Proof.
Denote by $c_{f}$ the coefficient of $x_{n+1}^{d}$ in $f$. Using the coordinate transformation from Lemma 4.2, we can assume that $h(c_{f})=h(f)$.
Let $M$ denote the candidate degree of $g$, and denote by $B[M]$ the set of homogeneous monomials in $x_{0},\ldots,x_{n+1}$ of degree $M$. Note that $|B[M]|=\binom{M+n+1}{n+1}$. Let $\Xi$ be a maximal subset of $X(b;P_{1},\ldots,P_{\ell})$ which is algebraically independent over $B[M]$ in the sense that applying the elements of $B[M]$ to $\Xi$ gives a matrix $A$ with linearly independent rows. So the columns of $A$ are parametrized by $B[M]$, while the rows are parametrized by $\Xi$. Let $r=|B[M]|$ and $s=|\Xi|$. If $g$ is a polynomial of degree $M$ vanishing on $\Xi$, then it actually vanishes on all of $X(b;P_{1},\ldots,P_{\ell})$. Assume that all degree $M$ polynomials vanishing on $X(b;P_{1},\ldots,P_{\ell})$ are divisible by $f$. We will obtain a contradiction when choosing $\ell$, the $\mu_{i}$, and $M$ as stated. For use below, we may already assume that $d^{2}=O_{n}(M)$.
The matrix $A$ describes a linear system whose solutions are the degree $M$ homogeneous polynomials vanishing on $\Xi$. By assumption, the kernel of $A$ is exactly $f\cdot B[M-d]$, and so we obtain that $s=|B[M]|-|B[M-d]|$. Let $\Delta$ be the greatest common divisor of all $s\times s$ minors of $A$. By Lemma 2.12, and the assumption on $c_{f}$, we have that
$h(\Delta)\leq h(\det(A\overline{A}^{T}))/2-(r-s)h(f).$
On the other hand, the divisibility from Lemma 4.1 above shows that
$h(\Delta)\geq\frac{n!^{1/n}n}{n+1}s^{1+1/n}\sum_{i=1}^{\ell}\left(\frac{1}{\mu_{i}^{1/n}}-O_{n}(s^{-1/n})\right).$
Write $\kappa=\sum_{i}\frac{1}{\mu_{i}^{1/n}}$. Putting these together, we have that
$\frac{n!^{1/n}n\kappa}{n+1}s^{1+1/n}-O_{n}(\ell s)\leq h(\det(A\overline{A}^{T}))/2-(r-s)h(f).$
The entries of $A$ are monomials of degree $M$ evaluated at points of height $<b$, and so $h(\det(A\overline{A}^{T}))/2\leq Ms(b-1)$. Dividing by $Ms$ yields that
$\frac{n!^{1/n}ns^{1/n}\kappa}{(n+1)M}-O_{n}(\ell/M)\leq b-1-\frac{r-s}{Ms}h(f).$
Now
$s$ $=|B[M]|-|B[M-d]|=\binom{M+n+1}{n+1}-\binom{M-d+n+1}{n+1}$
$=dM^{n}/n!+O_{n}(d^{2}M^{n-1}).$
So by using that $d^{2}/M=O_{n}(1)$, we have that $s^{1/n}/M=d^{1/n}/n!^{1/n}+O_{n}(d^{2}/M)$. Therefore the left-hand side can be replaced by
$\frac{d^{1/n}n}{n+1}\kappa-O_{n}\left(\frac{\ell d^{2}}{M}\right).$
For the right-hand side we have that
$\frac{r-s}{Ms}=\frac{1}{d(n+1)}+O_{n}({1/M}),$
and so the inequality becomes
$\frac{d^{1/n}n}{n+1}\kappa-O_{n}\left(\frac{\ell d^{2}}{M}\right)\leq b-1-\frac{h(f)}{d(n+1)}+O_{n}\left(\frac{h(f)}{M}\right).$
Rearranging, we obtain that
(2) $\kappa\leq\frac{(n+1)(b-1)}{nd^{1/n}}-\frac{h(f)}{nd^{1+1/n}}+O_{n}\left(\frac{h(f)}{d^{1/n}M}\right)+O_{n}\left(\frac{\ell d^{2-1/n}}{M}\right).$
We now take $M$ of size $O_{n}(d^{3}\ell+d^{2}b)$, where the implicit constant will be determined later. We distinguish two cases depending on $h(f)$. Assume first that $h(f)\leq 2(n+1)db$. By taking the implicit constant for $M$ large enough, we may arrange for the last two terms in Equation (2) to satisfy
$O_{n}\left(\frac{h(f)}{d^{1/n}M}\right)<\frac{1}{2nd^{1+1/n}},\quad O_{n}\left(\frac{\ell d^{2-1/n}}{M}\right)<\frac{1}{2nd^{1+1/n}}.$
We then obtain that
$\kappa<\frac{(n+1)(b-1)}{nd^{1/n}}-\frac{h(f)-1}{nd^{1+1/n}},$
as desired. Second, assume that $h(f)\geq 2(n+1)db$. By adjusting the implicit constant for $M$, we then have in Equation (2) that
$O_{n}\left(\frac{h(f)}{d^{1/n}M}\right)\leq\frac{h(f)}{2nd^{1+1/n}},\quad O_{n}\left(\frac{\ell d^{2-1/n}}{M}\right)<1/d,$
which leads to a contradiction as $\kappa$ is a positive number
$\kappa\leq\frac{(n+1)(b-1)}{nd^{1/n}}-\frac{2(n+1)db}{nd^{1+1/n}}+\frac{(n+1)db}{nd^{1+1/n}}+\frac{1}{d}<0.$
In conclusion, we obtain the desired contradiction by taking
$\kappa\geq\frac{(n+1)(b-1)}{nd^{1/n}}-\frac{h(f)-1}{nd^{1+1/n}}$
and $M$ of size $O_{n}(d^{1+1/n}b+d^{3}\ell)$.
Using a technique due to Ellenberg–Venkatesh [EV05], we obtain the following affine variant. Note that the proof in this case is much simpler than in e.g.  [CCDN20, Prop. 4.2.1], by the ubiquity of height one primes in ${\mathcal{O}}_{K}$.
Let $i\geq 0$. For $f\in{\mathcal{O}}_{K}[x_{1},\ldots,x_{n+1}]$, we denote by $f_{i}\in{\mathcal{O}}_{K}[x_{1},\ldots,x_{n+1}]$ its degree $i$ homogeneous part, which is a homogeneous polynomial.
Proposition 4.4 (Affine auxiliary polynomials).
Let $f\in{\mathcal{O}}_{K}[x_{1},\ldots,x_{n+1}]$ be primitive, absolutely irreducible of degree $d$, and let $X\subset{\mathbb{A}}^{n+1}_{K}$ be the affine variety cut out by $f$. Let $b$ be a positive integer and take positive integers $\mu_{1},\ldots,\mu_{\ell}$ for which
$\sum_{i=1}^{\ell}\frac{1}{\mu_{i}^{1/n}}\geq\frac{b-1}{d^{1/n}}-\frac{h(f_{d})-1}{nd^{1+1/n}}.$
Let $p_{1},\ldots,p_{\ell}$ be primes. Then if $P=(P_{1},\ldots,P_{\ell})\in\prod X_{p_{i}}({\mathbb{C}})$ is a tuple where each $P_{i}$ is of multiplicity $\mu_{i}$, then there exists a polynomial
$g:=g_{P}\in{\mathcal{O}}_{K}[x_{1},\ldots,x_{n+1}]$
coprime to $f$, of degree $O_{n}(d^{1+1/n}b+d^{3}\ell)$, and vanishing on
$X(b;P_{1},\ldots,P_{\ell})=\{x\in X(b):x\equiv P_{i}\bmod p_{i}\}.$
Proof.
By Corollary 2.6 there exists a point $a\in{\mathbb{C}}^{n+1}$ which does not lie in $X(1)$. By considering $f(x+a)$, we may assume that $f_{0}$ is non-zero. Take an element $H\in{\mathcal{O}}_{K}$ of height $b-1$ which does not divide $f_{0}$. Consider the primitive homogeneous polynomial
$F(x_{0},x_{1},\ldots,x_{n+1})=\sum_{i=0}^{d}H^{i}f_{i}x_{0}^{d-i}.$
We note that $F$ is absolutely irreducible because $x_{0}\nmid F$ and $F(H,x_{1},\ldots,x_{n+1})=H^{d}f(x_{1},\ldots,x_{n+1})$, and that $h(F)\geq d(b-1)+h(f_{d})$. Every point $(x_{1},\ldots,x_{n+1})$ on $X$ gives a point $(H:x_{1}:\ldots:x_{n+1})$ on $V(F)$. By applying Theorem 4.3 to $F$ we get the desired result. ∎
(4.3). Counting on curves
Using Theorem 4.3 and Proposition 4.4 we can quickly deduce results on counting on projective and affine curves.
Proof of Theorem D.
By Lemma 2.9 we may assume that $n=2$, i.e. $C$ is a planar curve. We may moreover assume that $C$ is geometrically integral, by Lemma 2.7. Put $\ell=\left\lfloor\frac{2(b-1)}{d}\right\rfloor+1$. Let $Z_{1},\ldots,Z_{k}$ be the irreducible components of $C(b)$ of maximal dimension. If $Z_{j}$ is completely contained in the singular locus $C^{\mathrm{sing}}({\mathbb{C}}(t))$, then $Z_{j}$ is a finite set and hence we are done. Otherwise we fix on every $Z_{j}$ a point $Q_{j}$ which is smooth as a point on $C$. Since being smooth and being geometrically integral are generic conditions, there exist non-zero primes $p_{1},\ldots,p_{\ell}\in{\mathcal{O}}_{K}$ such that each reduction $C_{p_{i}}$ is geometrically integral, and such that for every $j$ the point $Q_{j}$ reduces to a smooth point on every $C_{p_{i}}$. Let $Y\subset C(b)$ be the subset of points $P\in C(b)$ which reduce to a smooth point on every $C_{p_{i}}$, and note that $Q_{j}\in Y$ for every $j$. Then combining Theorem 4.3 with all $\mu_{i}=1$ and Bézout’s theorem shows that the morphism
$Y\to\prod_{i}C_{p_{i}}$
is finite-to-one onto its image, and has degree bounded by $O(d^{3}b)$. Now, for each irreducible component $Z_{j}$ the set of $Q\in Z_{j}$ which reduce to a smooth point on every $C_{p_{i}}$ form an open subset of $Z_{j}$. It follows that $Y$ is dense in every $Z_{j}$. The result now follows by noting that $Y$ has at most $O(d^{3}b)$ irreducible components of dimension $\ell=\dim(\prod_{i}C_{p_{i}})$. ∎
Proof of Theorem C.
The argument is identical to the previous corollary, but using Proposition 2.8 instead of Lemma 2.9, and Proposition 4.4 instead of Theorem 4.3. ∎
5. Geometric dimension growth
In this section we prove our main results on dimension growth. The projective case will be deduced from the affine case, which we will prove by induction on dimension. The base case for Theorem B is counting on affine surfaces.
For the affine case, we will prove a dimension growth result which is more general than what is needed for Theorem B, similar to [Ver24, CDH+23].
We embed ${\mathbb{A}}^{n}_{K}$ in ${\mathbb{P}}^{n}_{K}$ where the hyperplane at infinity $H_{\infty}$ is given by $x_{0}=0$. If $X\subset{\mathbb{A}}^{n}_{K}$ is an affine variety, we denote by $\overline{X}$ the Zariski closure of $X$ in ${\mathbb{P}}^{n}$. Let $X_{\infty}=\overline{X}\cap H_{\infty}$, which is a projective variety in ${\mathbb{P}}^{n-1}_{K}$. If $X$ is defined by a degree $d$ polynomial $f\in K[x_{1},\ldots,x_{n}]$, then $X_{\infty}$ is defined by $f_{d}=0$ in ${\mathbb{P}}^{n-1}_{K}$.
Theorem 5.1 (Dimension growth for affine varieties).
Let $X\subset{\mathbb{A}}^{n}_{K}$ be an irreducible affine variety of dimension $m$ and of degree $d\geq 6$ Assume that $X_{\infty}$ has at least one geometrically irreducible component of degree at least $3$. Then for all $b\geq 1$
$N_{\mathrm{aff}}(X,b)\leq(m-1)b+1,$
and the number of irreducible components of $X(b)$ of dimension $(m-1)b+1$ is at most $O(d^{7})$. For $b\geq 7$ this number of irreducible components is at most $O(d^{4})$.
Remark 5.2.
Note that the condition that $X$ is not cylindrical over a curve (abbreviated NCC) appearing in the main affine results of [Ver24, CDH+23] holds if $X_{\infty}$ has at least one geometrically irreducible component of degree at least $3$ as above. Indeed, a hypersurface $X=V(f)$ is NCC if it is not the preimage of a planar curve under a $K$-linear map. The condition on $X_{\infty}$ guarantees $f_{d}$ is not a form in two variables after any $K$-linear change of coordinates.
(5.1). Surfaces
We first count on affine surfaces. The goal of this section is to prove the following proposition:
Proposition 5.3.
Let $f\in{\mathcal{O}}_{K}[x,y,z]$ be an absolutely irreducible polynomial of degree $d\geq 6$, and let $X=V(f)\subset{\mathbb{A}}^{3}_{K}$ be the corresponding surface. Assume that $f_{d}$ has at least one absolutely irreducible factor of degree at least $3$. Then for $b\geq 1$ we have
$N_{\mathrm{aff}}(X,b)\leq b+1.$
For $b\geq 7$, the number of irreducible components of $X(b)$ of dimension $b+1$ is at most $O(d^{4})$, while for $b<7$ this number is $O(d^{7})$.
Moreover, if $f_{d}$ is absolutely irreducible and $U\subset X$ is the open subset obtained by removing all lines from $X$, then
$N_{\mathrm{aff}}(U,b)\leq\max\left\{\frac{b+1}{2},\left(\frac{3}{2\sqrt{d}}+\frac{1}{3}\right)(b-1)+\frac{5}{2}\right\}.$
For the proof, we will need a result on rulings of surfaces by curves of degree $1$ and $2$. First of all let us state what we mean by a ruling. By a conic in ${\mathbb{A}}^{3}_{K}$ we mean a curve of degree $2$. Note that every conic must lie on a plane.
Definition 5.4.
Let $X\subset{\mathbb{P}}_{K}^{3}$ be a projective surface.
  1. (1)
    We say that $X$ is ruled if there exist infinitely many distinct lines contained in $X$.
  2. (2)
    We say that $X$ ruled tangentially by conics if there exist infinitely many distinct conics contained in $X$ which are tangent to $H_{\infty}$.
In the above context, by referring to the ruling $\mathcal{R}$, we mean the respective sets of lines and conics satisfying the given conditions.
Note that in the above definition the conics may be degenerate, i.e. they can be a double line.
Lemma 5.5.
Let $X\subset{\mathbb{P}}_{K}^{3}$ be a projective, geometrically integral surface of degree $d\geq 2$ and suppose that $X$ is ruled or tangentially ruled by conics with a ruling $\mathcal{R}$. Assume that $X_{\infty}$ has at least one geometrically irreducible component $E$ of degree at least $3$. Then the following hold.
  1. (1)
    $E$ is the unique geometrically irreducible component of $X_{\infty}$ of degree at least $3$.
  2. (2)
    For every $x\in X$ there exists a $C_{x}\in{\mathcal{R}}$ containing $x$ which intersects $X_{\infty}$ in $E$.
  3. (3)
    For every $e\in E$ there are only finitely many $C\in\mathcal{R}$ passing through $e$.
  4. (4)
    There are only finitely many $C\in\mathcal{R}$ which do not intersect $E$.
Proof.
First of all we construct a projective $K$-variety $Z$ parameterizing conics in ${\mathbb{P}}^{3}$. Note that every conic in ${\mathbb{P}}^{3}$ is the intersection of a quadratic form and a hyperplane. Hyperplanes in ${\mathbb{P}}^{3}$ are naturally parameterized by the dual space $({\mathbb{P}}^{3})^{*}\cong{\mathbb{P}}^{3}$, say with coordinates $[q_{0}:q_{1}:q_{2}:q_{3}]$. Consider an affine chart $q_{i}\neq 0$ in $({\mathbb{P}}^{3})^{*}$ for some $0\leq i\leq 3$. We can then write $x_{i}=-\sum_{j\neq i}\frac{q_{j}}{q_{i}}x_{j}$, so that the intersection of a quadratic form with this hyperplane will be entirely determined by a quadratic form $Q$ in the $x_{j}$, $j\neq i$. The space of such quadratic forms is isomorphic to ${\mathbb{P}}^{5}$. We then get a scheme $Z_{i}\cong{\mathbb{A}}^{3}\times{\mathbb{P}}^{5}$. We construct $Z$ by gluing together the $X_{i}$ in the obvious way.
Note that $Z$ comes equipped with a natural map $Z\to{\mathbb{P}}^{3}$, projecting to the hyperplane containing the conic, and that $Z$ is a ${\mathbb{P}}^{5}$-bundle over ${\mathbb{P}}^{3}$, implying that $Z$ is complete. Note also that conics which are in fact double lines are not uniquely parameterized by $Z$, since they are contained in infinitely many hyperplanes.
Next, consider the subvariety $Z_{\infty}$ of $Z$ consisting of those conics tangent to $H_{\infty}$. This is clearly a closed condition on $Z$. Let $Y\subset Z_{\infty}$ be the subvariety consisting of conics which are contained in $X$. We argue that this is a closed subvariety of $Z$. Consider
=${\mathcal{I}}=\{(x,\gamma)\in{\mathbb{P}}^{3}\times Z_{\infty}:x\in\gamma\},$
which is clearly a closed subvariety of ${\mathbb{P}}^{3}\times Z_{\infty}$, and denote by $\pi_{1}:{\mathcal{I}}\to{\mathbb{P}}^{3}$ and $\pi_{2}:{\mathcal{I}}\to Z_{\infty}$ the coordinate projections. Since ${\mathbb{P}}^{3}$ is complete, we see that $\pi_{2}(\pi_{1}^{-1}(X))=Y$ is closed in $Z_{\infty}$, as desired.
Whether we have a ruling $\mathcal{R}$ by lines or by conics, we must have a geometrically irreducible component $D$ of $X_{\infty}$ such that infinitely many different curves $C\in\mathcal{R}$ intersect $D$. Consider the projective variety
$T:=\{(p,q,L)\in X\times D\times Y:p,q\in L\}.$
The image of the projection $\pi:T\to X$ must be a closed subset of $X$. By assumption, this image contains infinitely many distinct $C\in\mathcal{R}$. As $X$ is geometrically integral, we therefore have that $\pi(T)=X$. Hence, for every $p\in X$, there is a curve $C\in\mathcal{R}$ lying on $X$ passing through $D$.
In particular, through every point of $E\setminus D$ there is either a line or conic contained in $X_{\infty}$ intersecting $D$. As $E$ has degree at least $3$, and is geometrically irreducible, this would mean that $X_{\infty}$ has infinitely many irreducible components, a contradiction. Hence $D=E$ for every irreducible component $E$ of $X_{\infty}$ of degree at least $3$, proving (1), (2) and (4).
For (3), note that if this statement would be false, then there exists some $e\in E$ such that for every point $x\in X$ there exists a line (resp. conic) through $e$ and $x$ contained in $X$. But then a similar argument as above shows that $X_{\infty}$ has infinitely many irreducible components. ∎
We first study the contribution to counting points of bounded height from lines and conics on the surface. For lines, we have the following lemma.
Lemma 5.6.
Suppose $f\in\mathcal{O}_{K}[x,y,z]$ is absolutely irreducible of degree $d\geq 2$ and let $X=V(f)\subset{\mathbb{A}}^{3}_{K}$. Assume that $f_{d}$ has at least one absolutely irreducible factor of degree at least $3$. Let $b$ be a positive integer, and denote by $Y\subset X$ the union of all lines $\ell$ on $X$ such that $\ell(b)$ contains at least two distinct points of $X(b)$. Then $Y(b)$ is a constructible set over ${\mathbb{C}}$, we have
$\dim Y(b)\leq b+1,$
and the number of irreducible components of $Y(b)$ of dimension $b$ is $O(d^{4})$.
Proof.
If $L$ is a line in ${\mathbb{A}}^{3}_{K}$ then it is of the form
$L=\{a+\lambda v:\lambda\in K\},$
for some $a\in{\mathbb{A}}^{3}_{K}$ and $v=(v_{1},v_{2},v_{3})\in{\mathcal{O}}_{K}^{3}\setminus\{(0,0,0)\}$, where we may assume that the $v_{i}$ are coprime. This line is contained in $X$ if and only if for every $\lambda\in K$ we have that $a+\lambda v\in X$. By Corollary 2.6 applied to $L\cap X\subseteq L$, it is enough to require this condition for all $\lambda\in{\mathbb{C}}$. If $L$ contains at least two points on $X(b)$, then we may assume that $a\in X(b)$ and that $h(v)<b$. Consider the set
$\{(a,v)\in X(b)\times({\mathbb{A}}^{3}_{K}\setminus\{(0,0,0)\})(b):a+\lambda v\in X(b)\,\,\forall\lambda\in{\mathbb{C}}\}.$
By Chevalley’s theorem, this is a constructible set over ${\mathbb{C}}$, and its projection onto the first coordinate is exactly $Y(b)$ and hence constructible.
Now let $L$ be such a line in $Y$, with base-point $a\in X(b)$ and $v\in{\mathcal{O}}_{K}^{3}$ primitive. Then the set of integral points on $L$ is given exactly by
$L({\mathcal{O}}_{K})=\{a+\lambda v:\lambda\in{\mathcal{O}}_{K}\}$
and hence $N_{\mathrm{aff}}(L,b)=b-h(v)$ and $L(b)$ is irreducible.
If $v$ is a direction of a line $L$ on $Y(b)$, then by expanding $f(a+\lambda v)$ into powers of $\lambda$, we see that $f_{d}(v)=0$. As discussed in Remark 5.2, $X$ is NCC and so [Ver24, Lem. 4.7] shows that there can be at most $d^{2}$ lines contained in $X$ which are parallel to a given line $L$ on $X$. Combining these facts, the strategy is to bound the integral points on lines by counting possible directions for lines, which come from rational points on the curve $X_{\infty}\subset{\mathbb{P}}^{2}_{K}$ defined by $f_{d}=0$.
Let $E$ be a geometrically irreducible component of $X_{\infty}$ which rules $X$ as in Lemma 5.5 if it exists. Otherwise put $E=\varnothing$. Lemma 5.5 shows that $E$ is unique and of degree $e\geq 3$, if it exists.
We first count on lines not coming from $E$. Let $g\in K[x,y,z]$ be the homogeneous factor of $f_{d}$ obtained by removing $E$, i.e. the vanishing locus of $g$ is exactly the Zariski closure of $X_{\infty}\setminus E$. Then Lemma 5.5 shows that projecting the set
$\{(a,v)\in{\mathbb{A}}^{3}\times{\mathbb{P}}^{2}_{K}:g(v)=0\wedge\forall\lambda\in K\left(f(a+\lambda v)=0\,\right)\}$
onto its first coordinate yields a finite union of lines on $X$. Expanding $f(a+\lambda v)$ as a polynomial in $\lambda$ gives polynomials of degree at most $d$ in $a,v$. Therefore [Ver24, Lem. 4.6] implies that this set consists of at most $d^{4}$ lines.
All other lines on $X$ must come from the ruling and intersect $E$. Given an integer $k$ with $1\leq k\leq b$, denote by $Y_{k}\subset Y(b)$ the subvariety of points on a line $L$ whose direction $v$ satisfies $v\in E$ and $h(v)=k-1$. Recall that for each such line $L$ we have $\dim L(b)\leq b-k+1$. We will show that $\dim Y_{k}\leq b$ and count the number of irreducible components of $Y_{k}$. By Theorem D, as $e\geq 3$, we get that
$\dim E(k)\leq\left\lfloor\frac{2(k-1)}{e}\right\rfloor+1.$
In particular, for each $k\geq 1$, the dimension of $Y_{k}$ is at most
$\dim Y_{k}\leq\dim E(k)+(b-k+1)\leq\left\lfloor\frac{2(k-1)}{e}\right\rfloor+b-k+2\leq b+1.$
For $k\geq 2$ this shows that the dimension of $Y_{k}$ is at most $b$. We thus only have to consider $Y_{1}$ in counting irreducible components.
To count the number of irreducible components, note that $E(1)$ has at most $e$ irreducible components of dimension $1$, by Corollary 2.5. For a line $L$ on $X$ there are at most $d^{2}$ lines on $X$ parallel to it, and hence $Y_{1}$ has at most $ed^{2}$ irreducible components of dimension $b$. Putting everything together, we obtain that the number of irreducible components of $Y(b)$ of dimension $b+1$ is bounded by
$O\left(d^{4}+d^{2}e\right)\leq O(d^{4}).\qed$
For conics on a surface we have the following.
Lemma 5.7.
Suppose $f\in\mathcal{O}_{K}[x,y,z]$ is absolutely irreducible of degree $d\geq 3$ and let $X=V(f)\subset{\mathbb{A}}^{3}_{K}$. Assume that $f_{d}$ has an absolutely irreducible factor of degree at least $3$.
Let $b$ be a positive integer, and denote by $Y\subset X$ the union of all conics on $X$ which are tangent to the plane at infinity in a $K$-rational point. Then $Y(b)$ is a constructible set over ${\mathbb{C}}$ with
$\dim Y(b)\leq\frac{2b+1}{3}.$
If $f_{d}$ is absolutely irreducible and $d\geq 4$ then one may improve this to
$\dim Y(b)\leq\frac{b+1}{2}.$
Moreover, if $C$ is a conic lying on $X$ and not tangent to the plane at infinity, then $C(b)$ is at most one-dimensional for all $b$.
Proof.
It follows that $Y(b)$ is a variety over ${\mathbb{C}}$ in the same way as it did for ruling by lines. We will follow the strategy of Browning–Heath-Brown–Salberger [BHBS06] in Section 6.1 to prove the other claims.
Let $C$ be any non-degenerate conic contained in $X$. If $C(b)$ is of dimension 00, then we are done. Otherwise, $C$ is determined in ${\mathbb{P}}^{3}$ by a linear form $L(x_{0},x_{1},x_{2},x_{3})$ and a quadratic form $q(x_{0},x_{1},x_{2},x_{3})$, both with coefficients in ${\mathcal{O}}_{K}$.
If $C$ is not tangent to $H_{\infty}$, then the result follows by Proposition 1.1 and Proposition 2.8. So we may assume that $C$ is tangent to $H_{\infty}$. As $q$ has coefficients in ${\mathcal{O}}_{K}$, the conic $C$ meets $H_{\infty}$ at a point $y\in H_{\infty}$. We will show that $N_{\text{aff}}(C,b)\leq\frac{b-h(y)}{2}$. Our strategy will be to show that the height $b$ points on $C$ can be parameterized by finitely many quadratic polynomials of a certain form.
We may assume without loss of generality that $L(x)=\sum_{i}a_{i}x_{i}$ with $a_{3}\neq 0$ for some $a_{i}\in{\mathcal{O}}_{K}$. Hence we have on $C$ that
$x_{3}=a_{3}^{-1}(a_{0}x_{0}+a_{1}x_{1}+a_{2}x_{2}).$
By substituting this in the quadratic form $q$ defining $C$, we get a quadratic form $Q(x_{0},x_{1},x_{2})$ which, together with the equation for $x_{3}$ above, determines $C$. As $C$ is tangent to $H_{\infty}$, we have that $Q(0,x_{1},x_{2})=a(\alpha x_{1}+\beta x_{2})^{2}$, where we can assume that $\alpha$ and $\beta$ have no common factors. Altogether we get,
$Q(x_{0},x_{1},x_{2})=a(\alpha x_{1}+\beta x_{2})^{2}+bx_{0}x_{1}+cx_{0}x_{2}+dx_{0}^{2}.$
We will reparameterize such that $\alpha x_{1}+\beta x_{2}$ becomes a single variable; we will do this in a ${\mathcal{O}}_{K}$-invertible way.
As ${\mathcal{O}}_{K}$ is a PID, we can find elements $\gamma$ and $\delta$ such that $\alpha\delta-\beta\gamma=1$. Denote $y_{0}=x_{0}$, $y_{1}=\alpha x_{1}+\beta x_{2}$, $y_{2}=\gamma x_{1}+\delta x_{2}$. In terms of $y_{0}$, $y_{1}$, $y_{2}$ the quadratic form then becomes
$Q(y_{0},y_{1},y_{2})=ay_{1}^{2}+ey_{0}y_{1}+gy_{0}y_{2}+dy_{0}^{2}.$
Note that $g\neq 0$ as $C$ is non-degenerate. Now, using $Q(y_{0},y_{1},y_{2})=0$, we see that we need
$y_{2}=\frac{-ay_{1}^{2}-dy_{0}^{2}-ey_{0}y_{1}}{gy_{0}}$
if $y_{0}\neq 0$. Because we care about the affine coordinates on $C$, we will be assuming $y_{0}=1$. Putting $y_{0}=x_{0}=1$, we then find that we can calculate $y_{2}$ in terms of $y_{1}$; hence we can find $x_{1}$, $x_{2}$ and $x_{3}$ in terms of $y_{1}$. In particular, we have quadratic polynomials $q_{1},q_{2},q_{3}$ over $K$ such that $x_{1}=q_{1}(y_{1})$, $x_{2}=q_{2}(y_{1})$, $x_{3}=q_{3}(y_{1})$ is a parameterization of $C$.
Now, if $C$ contains no ${\mathcal{O}}_{K}$-points of height at most $b$, we are done. Otherwise, let $\delta$ be an ${\mathcal{O}}_{K}$-point on $C$ of height at most $b$, and take $y^{*}\in K$ such that $\delta=(q_{1}(y^{*}),q_{2}(y^{*}),q_{3}(y^{*}))$. Note that $y^{*}=\alpha\delta_{1}+\beta\delta_{2}$, showing that in fact $y^{*}\in{\mathcal{O}}_{K}$. We can then change coordinates to $u=y-y^{*}$. Abusing notation by putting $x_{i}=q_{i}(u)$, we obtain $\delta_{i}=q_{i}(0)$, and hence we may assume that the $q_{i}$ have constant term of height at most $b$ in ${\mathcal{O}}_{K}$.
Next, let $D$ be a common denominator for all the $q_{i}$. Suppose
$q_{i}(u)=\delta_{i}+\frac{1}{D}(B_{i}u+C_{i}u^{2}).$
We want to find those $u\in{\mathcal{O}}_{K}$ for which $q_{i}(u)$ is in ${\mathcal{O}}_{K}$. We will subdivide them according to their greatest common divisor with $D$, so fix a factor $\lambda$ of $D$ and fix some $u\in{\mathcal{O}}_{K}$ for which $q_{i}(u)\in{\mathcal{O}}_{K}$ and $\gcd(u,D)=\lambda$.
Write $u=\lambda z$, $D=\lambda D^{\prime}$. We get that
$q_{i}(u)-\delta_{i}=\frac{1}{D^{\prime}}(B_{i}z+\lambda C_{i}z^{2})\in{\mathcal{O}}_{K}$
for all $i$. By assumption, $z$ has no common factor with $D^{\prime}$, so we have $D^{\prime}\mid B_{i}+\lambda C_{i}z$. Let $G_{i}$ be the greatest common divisor of $\lambda C_{i}$ and $D^{\prime}$. Then we obtain $G_{i}\mid B_{i}$. Denote $\lambda C_{i}=G_{i}C_{i}^{\prime}$, $B_{i}=G_{i}B_{i}^{\prime}$ and $D^{\prime}=G_{i}D_{i}^{\prime\prime}$. Then we find that $z\equiv-C_{i}^{\prime-1}B^{\prime}_{i}\pmod{D_{i}^{\prime\prime}}$, which determines $z$ uniquely modulo the least common multiple $d^{\prime}$ of the $D_{i}^{\prime\prime}$. Denote by $Z\in{\mathcal{O}}_{K}$ the element of smallest degree such that $Z\equiv-C_{i}^{\prime-1}B^{\prime}_{i}\pmod{D_{i}^{\prime\prime}}$.
We can then substitute $z=Z+d^{\prime}v$, with $v\in{\mathcal{O}}_{K}$, to get representations $q_{i}(v)$ (in terms of $v$) which are contained in ${\mathcal{O}}_{K}$.
Because we can repeat this for every factor $\lambda$ of $D$ (which is each time the greatest common divisor with $u$), we find finitely many quadratic representations $(q_{1i}(v),q_{2i}(v),q_{3i}(v))$ whose ${\mathcal{O}}_{K}$-images together cover every ${\mathcal{O}}_{K}$-point of $C$.
We can almost conclude the proof of the claim now. Focus on one representation
$q_{i}(v)=A_{i}+B_{i}v+C_{i}v^{2}.$
If it contains a height at most $b$ point in its ${\mathcal{O}}_{K}$-image, we may assume that the $A_{i}$ have height at most $b$, otherwise we are done.
Suppose without loss of generality that $\deg C_{1}\geq\deg C_{2},\deg C_{3}$. Write $B_{1}=B_{1}^{\prime}+D_{1}C_{1}$, with $\deg B_{1}^{\prime}<\deg C_{1}$. Then it is clear that in order to have $\deg q_{1}(V)<b$, we need $\deg v(v+D_{1})<b-\deg C_{1}$. If $\deg D_{1}<\frac{b-\deg C_{1}}{2}$, we see that the only requirement is $\deg v<\frac{b-\deg(C_{1})}{2}$, which means that $v$ can take values in an irreducible space of dimension $\frac{b-\deg C_{1}}{2}$ over ${\mathbb{C}}$.
If $\deg D_{1}\geq\frac{b-\deg C_{1}}{2}$, we have to distinguish between two cases:
  • If $\deg v\neq\deg D_{1}$, then $\deg v<b-\deg C_{1}-\deg D_{1}\leq\frac{b-\deg C_{1}}{2}$.
  • If $\deg v=\deg D_{1}$, then $\deg(v+D_{1})<\frac{b-\deg C_{1}}{2}$.
In either case $v$ can only take values in an irreducible space of dimension at most $\frac{b-\deg C_{1}}{2}$ over ${\mathbb{C}}$. We are now done once we realize that $[0:C_{1}:C_{2}:C_{3}]=y$, and hence $h(y)\leq\deg C_{1}$.
Suppose now that there are only finitely many non-degenerate conics lying on $X$ which are tangent to $H_{\infty}$. By the above, in this case $\dim C(b)\leq\frac{b}{2}$, so in particular $\dim Y(b)\leq\frac{b}{2}$.
Finally, suppose that there are infinitely many non-degenerate conics lying on $X$ which are tangent to $H_{\infty}$. By Lemma 5.5, all but finitely many of these conics pass through $E$, which corresponds to the unique absolutely irreducible factor of $f_{d}$ of degree at least $3$. We can ignore these finitely many conics and assume all conics we consider pass through $E$.
For $y\in E$, let $Y_{y}$ be the subvariety of $Y(b)$ consisting of those conics intersecting $H_{\infty}$ in $y$. We know that $Y_{y}$ is the union of finitely many $C(b)$ by Lemma 5.5. Hence
$\dim Y_{y}\leq\frac{b-h(y)}{2}.$
Using Theorem D, we conclude that
$\dim Y(b)\leq\max_{1\leq k\leq b}\frac{b-k}{2}+\left\lfloor\frac{2(k-1)}{3}\right\rfloor+1\leq\frac{2b+1}{3}.$
Note that the latter is strictly smaller than $b$ unless $b=1$. If $f_{d}$ is absolutely irreducible and $d\geq 4$, then $E$ has degree $d$, so that
$\dim Y(b)\leq\max_{1\leq k\leq b}\frac{b-k}{2}+\left\lfloor\frac{2(k-1)}{d}\right\rfloor+1\leq\frac{b+1}{2}.\qed$
Lemma 5.8.
Let $X\subset{\mathbb{A}}_{K}^{3}$ be a geometrically integral surface. Let $U\subset X$ be the subset of smooth points $x\in X$ for which the intersection multiplicity of $X$ with the tangent space $T_{x}X$ at $x$ is at most $2$. Then $X\setminus U$ consists of $O(d^{4})$ irreducible components of dimension one.
Proof.
This follows from [BHBS06, Lem. 10, 11]. Indeed, note that the universal polynomials $\Phi_{i}(a_{e};x_{0},x_{1},x_{2},x_{3})$ constructed there are of degree at most $O(d^{4})$ in $x$. ∎
We need one more elementary lemma for the proof of Proposition 5.3.
Lemma 5.9.
Let $d\geq 6,e\geq 3$ and $b\geq 1$ be integers. Define $\ell=\lfloor(b-1)/\sqrt{d}\rfloor+1$ and $\ell^{\prime}=\lfloor(b-1)/e-\ell/(e-1)\rfloor+1$. Then
$12\ell+\ell^{\prime}\leq b+$ 1 .
Moreover, if $b\geq 7$ then $2\ell+\ell^{\prime}\leq b$.
Proof.
[…]
$2\ell+\ell^{\prime}\leq\frac{2(b-1)}{\sqrt{d}}+\frac{b-1}{e}+3-\frac{b-1}{\sqrt{d}(e-1)}-\varepsilon,$
for some small $0<\varepsilon<1/(\sqrt{d}(e-1))$. If $b\geq 20$, then we see that this quantity is indeed bounded as stated. So we may assume that $b<20$. Also, if $b=1$ or $b=2$ then a straightforward computation gives the desired conclusion, so we can assume that $3\leq b\leq 19$. Now, if $b-1<\sqrt{d}$ then $\ell=1$, and so we also obtain the desired conclusion. So we may assume that $\sqrt{d}\leq b-1\leq 18$. For fixed $b,d$ with $d\geq 6$ we note that $\ell^{\prime}$ goes to $1$ as $e$ goes to $\infty$. Hence for each $b,d$ there are only finitely many values of $e$ to check. Going through all finitely many remaining cases using a computer program yields the proof. ∎
We have all of the ingredients for Proposition 5.3.
Proof of Proposition 5.3.
By Lemma 2.7 we may assume that $X$ is geometrically integral. Let $Y_{1}\subset X(b)$ consist of all integral points of height at most $b$ on the union of all lines on $X$ which contain at least two points of $X(b)$. Similarly, let $Y_{2}\subset X(b)$ consist of all integral points of height at most $b$ on $X$ on conics on $X$ which are tangent to the plane at infinity. By Lemmas 5.6 and 5.7 we have
$\dim Y_{1}\leq b+1,\dim Y_{2}\leq\frac{2b+1}{3},$
and that $Y_{1}$ has at most $O(d^{4})$ irreducible components of dimension $b+1$. Let $Z\subset X$ consist of all points $x$ on $X$ which are either singular on $X$, or for which the tangent plane section has intersection multiplicity strictly greater than $2$ at $x$. Then Lemma 5.8 and the Schwartz–Zippel bound from Proposition 2.4 show that
$N_{\mathrm{aff}}(Z,b)\leq b,$
and the number of irreducible components of $Z(b)$ of dimension $b$ is bounded by $O(d^{4})$.
Let $(Z_{j})_{j}$ be the irreducible components of $X(b)$ of maximal dimension. If $Z_{j}$ is contained in $Z(K)$ then $Z_{j}$ is in fact contained in $Z(b)$. Hence by the reasoning above we have $\dim Z_{j}\leq b$. Otherwise, we take a point $Q_{j}$ on $Z_{j}$ which is not in $Z$, i.e. it is smooth on $X$ and has intersection multiplicity at most $2$ with the tangent plane section. Put $\ell=\lfloor(b-1)/\sqrt{d}\rfloor+1$, which will be used to apply the determinant method later. Then there exist non-zero primes $p_{1},\ldots,p_{\ell}\in{\mathcal{O}}_{K}$ such that $X_{p_{i}}$ is geometrically integral, such that the reduction of $Q_{j}$ is still smooth on $X_{p_{i}}$, and such that the reduction still has intersection multiplicity at most $2$ with the tangent plane section at $Q_{j}\bmod p_{i}$. Indeed, all of these are generic conditions.
Let $Y\subset X(b)$ be the subset of points $P\in X(b)$ which reduce to a point which is smooth and has intersection multiplicity at most $2$ with the tangent plane section in every $X_{p_{i}}$. Consider the morphism
$\varphi:Y\to\prod_{i=1}^{\ell}X_{p_{i}}$
arising from reductions modulo $p_{i}$, $1\leq i\leq\ell$. Fix a point $P=(P_{1},\ldots,P_{\ell})$ in the image of this map. By Proposition 4.4, the preimage $\varphi^{-1}(P)\cap Y$ is contained in a set of the form $C(b)$ for a curve $C\subset X$ of degree at most $O(d^{7/2}b)$. In fact, $\varphi^{-1}(P)\cap Y$ is clearly even contained in $C(b;P_{1},\ldots,P_{\ell})$, consisting of those points $P\in C(b)$ such that $P\equiv P_{i}\bmod p_{i}$. Our aim is to bound the dimension of this object.
Let $D$ be an irreducible component of $C$, say of degree $e$. We may assume that $D$ is geometrically integral by Lemma 2.7. We will bound the dimension of $D(b;P_{1},\ldots,P_{\ell})$. If $D(b;P_{1},\ldots,P_{\ell})$ contains at most one point, then there is nothing to prove. So assume that $D(b;P_{1},\ldots,P_{\ell})$ consists of at least two points. If $e=1$ then $D(b;P_{1},\ldots,P_{\ell})$ is contained in $Y_{1}$. If $e=2$ then $D(b;P_{1},\ldots,P_{\ell})$ is either contained in $Y_{2}$, or the final statement of Lemma 5.7 applies. We may disregard either case and assume that $e\geq 3$.
Note that $1\leq\operatorname{mult}_{P_{i}}(D_{p_{i}})\leq e$. If this multiplicity were equal to $e$, then $D_{p_{i}}$ would be a cone, and hence $D_{p_{i}}$ would be contained in the intersection of $X_{p_{i}}$ with the tangent plane $T_{P_{i}}X_{p_{i}}$. But this contradicts the fact that $P_{i}$ has intersection multiplicity at most $2<e$ on the tangent plane section. Hence $\operatorname{mult}_{P_{i}}(D_{p_{i}})<e$ for every $i$. Define the integer $\ell^{\prime}$ as
(3) $\ell^{\prime}=\ell^{\prime}(e)=\left\lfloor\frac{b-1}{e}-\frac{\ell}{e-1}\right\rfloor+1.$
We now claim that
$\dim D(b;P_{1},\ldots,P_{\ell})\leq\ell^{\prime}.$
To prove the claim, we reason as in the proof of Theorem D. Let $(D_{j})_{j}$ be the irreducible components of $D(b;P_{1},\ldots,P_{\ell})$ of maximal dimension. If $D_{j}$ is contained in $D^{\mathrm{sing}}(K)$, then $D_{j}$ is finite, and this certainly satisfies the desired dimension bound. Otherwise, fix a point $Q_{j}^{\prime}$ on $D_{j}$ which is smooth on $D$. Take non-zero primes $q_{1},\ldots,q_{\ell^{\prime}}\in{\mathcal{O}}_{K}$ for which $Q_{j}^{\prime}$ remains smooth on the reduction $D_{q_{j}}$, and for which $D_{q_{j}}$ is geometrically integral. Let $D^{\prime}\subset D(b;P_{1},\ldots,P_{\ell})$ consist of all points on $D(b;P_{1},\ldots,P_{\ell})$ which reduce to a smooth point on every $D_{q_{j}}$. Then Proposition 4.4 shows that the map
$D^{\prime}\to\prod_{j=1}^{\ell^{\prime}}D_{q_{j}}$
is finite-to-one onto its image, of degree at most $O(e^{2}b+e^{3}\ell^{\prime})$. Indeed, let $Q_{1}^{\prime\prime},\ldots,Q_{\ell^{\prime}}^{\prime\prime}$ be smooth points on $D_{q_{1}},\ldots,D_{q_{\ell}^{\prime}}$. Since
$\ell^{\prime}+\sum_{i=1}^{\ell}\frac{1}{\operatorname{mult}_{P_{i}}(D_{p_{i}})}\geq\frac{b-1}{e},$
Proposition 4.4 provides us with an auxiliary curve of degree at most $O(e^{2}b+e^{3}\ell^{\prime})$ containing $D(b;P_{1},\ldots,P_{\ell},Q_{1}^{\prime\prime},\ldots,Q_{\ell^{\prime}}^{\prime\prime})$ but not containing $D$.
As in the proof of Theorem D, we then obtain that $D^{\prime}$ is dense in every $D_{j}$ of maximal dimension, and so we conclude that $D(b;P_{1},\ldots,P_{\ell})$ has dimension at most $\ell^{\prime}=\ell^{\prime}(e)$ as in Equation (3).
To conclude the proof, we have shown that every irreducible component of a fibre of the map $\varphi$ which is not contained in $Y_{1}$ or $Y_{2}$ has dimension bounded by
$\max_{e\geq 3}\ell^{\prime}(e).$
Now $\prod_{i=1}^{\ell}X_{p_{i}}$ has dimension $2\ell$, and so the union of all irreducible components of fibres of $\varphi$ which are not contained in $Y_{1}$ or $Y_{2}$ has dimension bounded by
$2\ell+\max_{e\geq 3}\ell^{\prime}(e).$
By Lemma 5.9 this quantity is always bounded by $b+1$. So we conclude that $X(b)$ has dimension bounded by $b+1$.
For the number of irreducible components, Lemma 5.9 shows that for $b\geq 7$ only irreducible components from $Y_{1}$ or $Y_{2}$ can contribute. For these, Lemma 5.6 show that this number is at most $O(d^{4})$. When $b<7$, Proposition 4.4 shows that the map $\varphi:Y\to\prod_{i=1}^{\ell}X_{p_{i}}$ has degree at most $O(d^{7})$, and so $X(b)$ has $O(d^{7})$ irreducible components of dimension $b+1$.
For the moreover part of the statement, let $U\subset X$ be the open subset obtained by removing all lines from $X$. Then in the above argument we do not have to consider $Y_{1}$. For $Y_{2}$, we have by Lemma 5.7 that
$\dim Y_{2}(b)\leq\frac{b+1}{2}.$
The set $Z$ is a one-dimensional algebraic set. Let $Z^{\prime}\subset Z$ be obtained by removing all lines from $Z$, so that $U\cap Z\subset Z^{\prime}$. Then by Theorem C we have that
$\dim Z^{\prime}(b)\leq\left\lceil\frac{b}{2}\right\rceil.$
Finally, we have argued above that all components of $X(b)$ which do not come from $Y_{1},Y_{2}$ or $Z$ have dimension at most
$2\ell+\max_{e\geq 3}\ell^{\prime}(e)\leq\left(\frac{3}{2\sqrt{d}}+\frac{1}{3}\right)(b-1)+\frac{5}{2}.$
Therefore we conclude that $U(b)$ has dimension at most
$\max\left\{\frac{b+1}{2},\left(\frac{3}{2\sqrt{d}}+\frac{1}{3}\right)(b-1)+\frac{5}{2}\right\}.\qed$
(5.2). Higher-dimensional varieties
In this section we develop the necessary slicing arguments to deduce our main results.
Proposition 5.10.
Let $n\geq 3$ and let $X\subset{\mathbb{P}}^{n}_{K}$ be a geometrically integral hypersurface. Let $Z\subset({\mathbb{P}}^{n}_{K})^{*}$ consist of all hyperplanes $H$ for which $X\cap H$ is not geometrically integral. Then $\dim Z(1)<n$.
In particular, note that this implies that there are infinitely many hyperplanes $H\subset{\mathbb{P}}^{n}$ defined over ${\mathbb{C}}$ for which $X\cap H$ is geometrically integral.
Proof.
By Bertini’s Thereom, the variety of linear forms $\ell_{\alpha}$ over $K$ for which $f|_{\{\ell_{\alpha}=0\}}$ is reducible is a Zariski closed set strictly contained in the space of all linear forms on ${\mathbb{A}}^{n}_{K}$. We are now done by Corollary 2.6. ∎
Lemma 5.11.
Let $n\geq 4$, let $X\subset{\mathbb{A}}^{n}_{K}$ be a geometrically integral hypersurface and assume that $X_{\infty}$ has a geometrically irreducible component of degree at least $3$. Let $Z\subset({\mathbb{P}}^{n}_{K})^{*}$ consist of all hyperplanes $H$ for which $X\cap H$ is not geometrically integral, or for which $(X\cap H)_{\infty}$ has no geometrically irreducible component of degree at least $3$. Then $\dim Z(1)<n$.
Again, this implies the existence of infinitely many hyperplanes $H\subset{\mathbb{A}}^{n}$ defined over ${\mathbb{C}}$ for which $X\cap H$ is geometrically integral and has a geometrically irreducible component of degree at least $3$ at infinity.
Proof.
This follows by applying the previous proposition to $X$ and to the geometrically irreducible component of $X_{\infty}$ of degree at least $3$. ∎
(5.3). Proofs of main results
In this section we finish up and prove our main results.
Lemma 5.12.
Let $n\geq 3$, let $f\in K[x_{1},\ldots,x_{n}]$ be absolutely irreducible of degree $d$, and denote $X=V(f)\subset{\mathbb{A}}^{n}_{K}$. Assume that $f_{d}$ is not a power of a linear form. Let $Z\subset({\mathbb{P}}^{n}_{K})^{*}$ consist of all hyperplanes for which $(X\cap H)_{\infty}$ is the vanishing locus of a power of a linear form. Then $\dim Z(1)<n$.
Proof.
Let $D_{1},\ldots,D_{k}\subset H_{\infty}={\mathbb{P}}^{n-1}_{K}$ be the geometrically irreducible components of $X_{\infty}$. Then either $k>1$ or $k=1$ and $D=D_{1}$ is not linear.
Assume first that $k>1$, so that $D_{1}\neq D_{2}$. The set $Y\subset({\mathbb{P}}^{n-1}_{K})^{*}$ consisting of all hyperplanes $H^{\prime}\subset{\mathbb{P}}^{n-1}_{K}$ for which $H^{\prime}\cap D_{1}=H^{\prime}\cap D_{2}$ is a proper closed subvariety of $({\mathbb{P}}^{n-1}_{K})^{*}$. Hence so is the set $Z^{\prime}\subset({\mathbb{P}}^{n}_{K})^{*}$ of hyperplanes $H\subset{\mathbb{P}}^{n}_{K}$ for which $H\cap H_{\infty}\in Y$. But $Z$ is contained in $Z^{\prime}$, so we conclude by Corollary 2.5.
Now assume that $k=1$ and write $D=D_{1}$. Then $D$ is not linear. If $n\geq 4$ then Bertini’s theorem implies that the set $Y\subset({\mathbb{P}}^{n-1}_{K})^{*}$ consisting of all hyperplanes $H^{\prime}\subset{\mathbb{P}}^{n-1}_{K}$ for which $H^{\prime}\cap D$ is not geometrically integral is a proper closed subvariety of $({\mathbb{P}}^{n-1}_{K})^{*}$. But if $H\in Z$ then $H\cap H_{\infty}\in Y$ and so the result follows from Corollary 2.5. If $n=3$ then $D$ is a curve in ${\mathbb{P}}^{2}_{K}$ of degree at least $2$. Hence, if $H^{\prime}\subset{\mathbb{P}}^{2}_{K}$ is a generic line, then $H^{\prime}\cap D$ consists of $\deg D$ many isolated points. These isolated points cannot be the vanishing locus of a linear polynomial, from which the result follows. ∎
For Theorem A we require the following.
Theorem 5.13.
Let $X\subset{\mathbb{A}}^{n}_{K}$ be an irreducible hypersurface defined by $X=V(f)$ for $f\in K[x_{1},\ldots,x_{n}]$. Assume that $f$ is of degree $d\geq 2$, and that $f_{d}$ is not a power of a linear form. Then
$N_{\operatorname{aff}}(X,b)\leq(n-2)b+1.$
Proof.
If $X$ is irreducible but not geometrically irreducible then the result follows from Lemma 2.7. So we may assume that $X$ is geometrically irreducible.
We induct on $n$, where the case $n=2$ follows from Proposition 1.1. So assume that $n\geq 3$. By Lemma 5.12, there exists a hyperplane $H\subset{\mathbb{A}}^{n}_{K}$ defined over ${\mathbb{C}}$ such that $X\cap H$ is geometrically integral, and such that $(X\cap H)_{\infty}$ is not the vanishing locus of a power of a linear form. Without loss of generality, we may assume $H$ is defined by
$x_{n}=a_{1}x_{1}+\ldots a_{n-1}x_{n-1}.$
For $c\in K$, denote by $H_{c}$ the hyperplane defined by $x_{n}=a_{1}x_{1}+\ldots a_{n-1}x_{n-1}+c$. There are only finitely many $c\in{\mathbb{A}}^{1}_{K}$ for which $X\cap H_{c}$ is either not geometrically irreducible, or $(X\cap H_{c})_{\infty}$ is a power of a linear form. For these values of $c$, we use the Schwartz–Zippel bounds to obtain $N_{\operatorname{aff}}(X\cap H_{c},b)\leq(n-2)b$. For the other values of $c$, by induction we have that $N_{\operatorname{aff}}(X\cap H_{c},b)\leq(n-3)b+1$. Since $X(b)$ is the union of the $(X\cap H_{c})(b)$ for $a$ running over ${\mathbb{A}}^{1}_{K}(b)$, we have that $X(b)$ is of dimension at most $(n-2)b+1$. ∎
Proof of Theorem A.
Let $X\subset{\mathbb{P}}^{n}_{K}$ be an irreducible projective hypersurface of degree $d\geq 2$. Let $Y\subset{\mathbb{A}}^{n+1}_{K}$ be the affine cone over $X$. Then the conditions from Theorem 5.13 are satisfied, and so we immediately obtain that
$N(X,b)\leq N_{\operatorname{aff}}(Y,b)-1\leq(n-1)b.\qed$
Proof of Theorem 5.1.
Let $X\subset{\mathbb{A}}^{n}_{K}$ be an irreducible affine variety of degree $d\geq 6$ and of dimension $m\geq 2$. Assume that $X_{\infty}$ has a geometrically irreducible component of degree at least $3$. By Lemma 2.7 and the Schwartz–Zippel bound, we may assume that $X$ is geometrically integral. Using Lemma 2.8 we may assume that $X$ is a hypersurface in ${\mathbb{A}}^{n}_{K}$.
We now induct on $n$, where the base case $n=3$ is Proposition 5.3. So assume that $n\geq 4$. By Lemma 5.11 there exists a hypersurface $H$ defined over ${\mathbb{C}}$ such that $X\cap H$ is geometrically integral and such that $(X\cap H)_{\infty}$ has a geometrically irreducible component of degree at least $3$. Without loss of generality, $H$ is given by $x_{n}=a_{1}x_{1}+\ldots+a_{n-1}x_{n-1}+a_{n}$ for some $a_{i}\in{\mathbb{C}}$. For $c\in{\mathcal{O}}_{K}$ let $H_{c}$ be the hyperplane given by $x_{n}=c+\sum_{i=1}^{n-1}a_{i}x_{i}$. Then the set of $c\in K$ for which $X\cap H_{c}$ is not geometrically integral, or for which $(X\cap H_{c})_{\infty}$ has no geometrically irreducible component of degree at least $3$, is finite. Consider the map $g:{\mathbb{A}}^{n}_{K}\to{\mathbb{A}}^{1}_{K}$ defined by
$g(x)=x_{n}-\sum_{i=1}^{n-1}a_{i}x_{i},$
and note that if $x\in H_{c}$ then $g(x)=c$, and that since $H$ is defined over ${\mathbb{C}}$ we have $h(g(x))\leq h(x)$. Now, the map $f:X(b)\to{\mathbb{A}}^{1}_{K}(b)$ induced by $g$ is algebraic, and by induction all but finitely many fibres have dimension bounded by $(m-2)b+1$. The remaining finitely many bad fibres have dimension bounded by $(m-1)b$ by the Schwartz–Zippel bound. So we conclude that the dimension of $X(b)$ is bounded by $(m-1)b+1$. Moreover, by induction the number of irreducible components of $X(b)$ of dimension $(m-1)b+1$ is at most $O(d^{4})$ if $b\geq 7$ and $O(d^{7})$ otherwise. ∎
Proof of Theorem B.
We may assume that $X$ is geometrically integral. Let $Y\subset{\mathbb{A}}^{n+1}_{K}$ be the affine cone over $X$. Then $Y$ is geometrically integral since $X$ is. Moreover, $Y_{\infty}$ is equal to $X$, and in particular has a geometrically irreducible component of degree at least $3$. Applying Theorem 5.1 then gives that
$N(X,b)=N_{\mathrm{aff}}(Y,b)-1\leq mb.$
The bound on the number of irreducible components of suitable dimension similarly follows. ∎
Proof of Theorem E.
We may assume that $X$ is geometrically irreducible, and recall that $U$ is obtained from $X$ by removing all lines. Let $Y\subset{\mathbb{A}}^{4}_{K}$ be the affine cone over $U$, so that
$N(U,b)\leq N_{\mathrm{aff}}(Y,b)-1.$
Then $Y$ is a geometrically integral quasi-affine $3$-fold, and $Y_{\infty}$ is also geometrically integral. By Lemma 5.11, there exists a hyperplane $H\subset{\mathbb{A}}^{4}_{\mathbb{C}}$ defined over ${\mathbb{C}}$ for which $Y\cap H$ and $(Y\cap H)_{\infty}$ are both geometrically integral. Let $H$ be given by $a_{1}x_{1}+\ldots+a_{n}x_{n}=a_{0}$ for some $a_{i},a_{0}\in{\mathbb{C}}$. For $c\in{\mathbb{C}}(t)$ let $H_{c}\subset{\mathbb{A}}^{4}_{K}$ be the hyperplane given by $a_{1}x_{1}+\ldots+a_{n}x_{n}=c$. Now $(Y\cap H_{c})_{\infty}=(Y\cap H)_{\infty}$ is geometrically integral, and so for every $c\in{\mathbb{C}}(t)$ we have that $Y\cap H_{c}$ is geometrically integral.
Take any $c\in{\mathbb{C}}(t)\setminus\{0\}$, so that 00 does not lie in $H_{c}$. If $\ell$ is a line on $H_{c}\cap Y$, then because $Y$ is a cone with cone point 00 the plane containing $\ell$ through 00 would be contained in $Y$. But this would yield a line on $U$, which by construction cannot exist. Hence $H_{c}\cap Y$ contains no lines and we conclude from Proposition 5.3 that
$N_{\mathrm{aff}}(H_{c}\cap Y,b)\leq\max\left\{\frac{b+1}{2},\left(\frac{3}{2\sqrt{d}}+\frac{1}{3}\right)(b-1)+\frac{5}{2}\right\}.$
For $H_{0}\cap Y$ we have by Proposition 5.3 that $N_{\mathrm{aff}}(H_{0}\cap Y,b)\leq b+2$. By fibering over $c\in{\mathbb{A}}^{1}_{K}(b)$, we conclude as desired that
$N_{\mathrm{aff}}(Y,b)\leq\left(1+\max\left\{\frac{1}{2},\frac{3}{2\sqrt{d}}+\frac{1}{3}\right\}\right)b+2.\qed$